### v1

# tableReferences for Table S1

Abstract
Generative models are invaluable in many fields of science because of their ability to capture high-dimensional and complicated distributions, such as photo-realistic images, protein structures, and connectomes. How do we evaluate the samples these models generate? This work aims to provide an accessible entry point to understanding popular notions of statistical distances, requiring only foundational knowledge in mathematics and statistics. We focus on four commonly used notions of statistical distances representing different methodologies: Using low-dimensional projections (Sliced-Wasserstein; SW), obtaining a distance using classifiers (Classifier Two-Sample Tests; C2ST), using embeddings through kernels (Maximum Mean Discrepancy; MMD), or neural networks (Fréchet Inception Distance; FID). We highlight the intuition behind each distance and explain their merits, scalability, complexity, and pitfalls. To demonstrate how these distances are used in practice, we evaluate generative models from different scientific domains, namely a model of decision making and a model generating medical images. We showcase that distinct distances can give different results on similar data. Through this guide, we aim to help researchers to use, interpret, and evaluate statistical distances for generative models in science.
footnotetext: Authors listed in alphabetical order.
1 Introduction
Generative models that produce samples of complex, high-dimensional data, have recently come to the forefront of public awareness due to their utility in a variety of scientific, clinical, engineering, and commercial domains (Bond-Taylor et al., 2021). Prominent examples include StableDiffusion and DALL-E for generating photo-realistic images (Rombach et al., 2022), WaveNet (Oord et al., 2016) for audio synthesis, and Generative Pre-trained Transformer (GPT;  Radford et al. 2018; 2019, Brown et al. 2020) for text generation. Besides this new wave of generative models, different scientific disciplines have a long history of building data generating models which capture specific processes. For example in neuroscience, the occurrence of action potentials is modeled at all different levels of detail (e.g. single neuron voltage dynamics (Hodgkin & Huxley, 1952) or at a phenomenological level (Pillow et al., 2008)), whereas in e.g. astrophysics there exist various models to simulate galaxy formation (Somerville & Davé, 2015). Along generating novel synthetic samples, generative models can be leveraged for specific tasks, such as sample generation conditioned on class labels (e.g., diseased vs. healthy brain scans, molecules that can or cannot be synthesized; Urbina et al. 2022, class-conditional image generation; van den Oord et al. 2016, Dockhorn et al. 2022, forecasting future states of a dynamical system; Durstewitz et al. 2023, Jacobs et al. 2023, Brenner et al. 2022), data imputation (e.g., Vetter et al. 2023, Lugmayr et al. 2022), data augmentation for downstream tasks (Rommel et al., 2022), and many more (see also Table S1).
These powerful capabilities are enabled by the premise that generative models accurately learn a high-dimensional distribution from which we assume our dataset was sampled. The dimensions can correspond to anything from individual pixels or graphs to arbitrary features of physical or abstract objects. When aiming to build generative models that better capture the true underlying data distribution, we need to answer a key question: How accurately does our generative model mimic the true data distribution?
Manual inspection of generated samples can be a good first check, e.g., in image or audio generation, where we can directly assess the visual likeness or sound quality of the samples (Gerhard et al., 2013, Vallez et al., 2022, Jayasumana et al., 2023). In general, however, we would like to have quantitative distances to measure the similarity of distributions, for instance to compare and benchmark different models. Many measures have been proposed that provide a quantitative way to assess the similarity of two distributions based on various aspects of their moments, spread, central tendency, and overall probability density (Fig. 1). Some of these measures require likelihood evaluations, as is possible with generative models such as Gaussian Mixture Models, Normalizing Flows, Variational Autoencoders, autoregressive models or diffusion models (Bishop, 2006, Papamakarios et al., 2021, Box et al., 2015, Kingma & Welling, 2014, Yenduri et al., 2023, Ho et al., 2020, Song et al., 2021). However, many contemporary machine learning models (e.g., Generative Adversarial Networks and Energy-Based Models; Goodfellow et al. 2014, Rezende et al. 2014, Hinton et al. 2006) and scientific simulators (e.g., single neuron voltage dynamics; Hodgkin & Huxley 1952) only define the likelihood implicitly, i.e. we can not explicitly evaluate their likelihood. In this work we focus on distances that can be applied to both classes of models, i.e. distances that can be computed only based on generated samples.
Figure 1: The need for statistical distances in scientific generative modeling. (a) An example target distribution, $p_{true}(x)$, and two learned distributions ($p_{1}(x)$ and $p_{2}(x)$) of different models trained to capture $p_{true}(x)$. All three distributions share the same mean and marginal variances, despite having distinct shapes. However, an appropriate sample-based distribution distance $D$ can determine that $p_{2}(x)$ is more similar to $p_{true}(x)$. (b) Scientific applications often require evaluating high-dimensional distributions, such as distributions for images or tabular data. In this example, each point represents an X-ray image, where each dimension is one pixel.
Here, we provide a guide to understanding commonly used sample-based statistical distances. Note that with distance, we do not necessarily refer to a distance metric in the mathematical sense (i.e., satisfying symmetry and the triangle inequality) but to a general measure of dissimilarity between two distributions. Our goal is not to provide a comprehensive review of statistical distances, as there are already a number of excellent resources for that purpose, especially in specific domains of application (Borji, 2019, Xu et al., 2018, Lopez-Paz & Oquab, 2016, Lueckmann et al., 2021). We refer readers to those works for a deeper dive into mathematical properties and empirical comparisons  (Cox et al., 1984, Gibbs & Su, 2002, Basseville, 2013, Cai & Lim, 2022, Muandet et al., 2017, Theis et al., 2016, Betzalel et al., 2022). In this guide, we instead focus on four commonly applied sample-based distances in the machine learning literature for evaluating eventually high-dimensional generative models. They represent different methodologies for defining statistical distance: Using low-dimensional projections (Sliced-Wasserstein; SW), obtaining a distance using classifiers (Classifier Two-Sample Tests; C2ST), using embeddings through kernels (Maximum Mean Discrepancy; MMD) or neural networks (Fréchet Inception Distance; FID). We aim to provide an intuition for how and when to apply these distances, and to build a solid foundation for navigating the extensive literature on statistical distances.
Towards this goal, we provide for each of the four distances an intuitive and graphical explanation (Section 2). We then perform a systematic evaluation of their robustness as a function of dataset size, data dimensionality, and other factors, such as data multimodality (Section 3). Finally, in Section 4, we demonstrate how these distances can be applied to compare generative models in different scientific domains: We evaluate low dimensional models of decision making in behavioral neuroscience and generative models of medical X-ray images. We show the importance of using multiple complementary distances, as distinct distances can give different results when comparing the same sets of samples. By presenting these resources, we aim to empower researchers to choose, implement, and evaluate the use and outcomes of statistical distances for generative models in science.
2 Sample-based statistical distances
In this section, we provide an overview of four classes of sample-based statistical distances commonly used in machine learning literature. Each class takes a different approach to overcoming the challenges inherent in comparing samples from high-dimensional and complex distributions. Throughout the section, we assume that we want to evaluate the distance between two datasets of samples, denoted as $\{x_{1},x_{2},\ldots,x_{n}\}\sim p_{1}(x)$ and $\{y_{1},y_{2},\ldots,y_{m}\}\sim p_{2}(y)$, where $p_{1}(x)$ and $p_{2}(y)$ are two probability distributions. These can be either two generative models, or a generative model and the underlying distribution of the observed data.
2.1 Slicing-based: Sliced-Wasserstein (SW) distance
Figure 2: Schematic for the Sliced-Wasserstein distance. (a) Samples from two two-dimensional distributions along with example slices. The “slicing” is done by sampling random directions from the unit sphere and projecting the samples from the higher-dimensional distribution onto that direction. (b) One-dimensional projections of the two distributions corresponding to the two random slices in (a). For each pair of projections, the empirical Wasserstein distance is computed. Unlike in higher dimensions, this can be done efficiently for one-dimensional distributions.
Computing distance between distributions suffers from the curse of dimensionality, where the computational cost of computing the distance increases very rapidly as the dimensionality of the data increases. This problem is especially restricting when the distance is used as part of a loss function in optimization problems, since in this case it needs to be evaluated many times. This has prompted the notion of “sliced” distances, which have become increasingly popular in recent years (Kolouri et al., 2019, Nadjahi et al., 2020, Goldfeld & Greenewald, 2021). The main idea behind slicing is that for many existing statistical distances we can efficiently evaluate the distance in low-dimensional spaces, especially in one dimension. Therefore, the “slices” are typically one-dimensional lines through the data space (Fig. 2a). All data points from each distribution are projected onto this line by finding their nearest point on the line, giving a one-dimensional distribution of projected data points (Fig. 2b). The distance measure of interest between the resulting one-dimensional distributions can then be computed efficiently. However, computing the random projection could lead to an unreliable measure of distance: Distinct distributions can produce the same one-dimensional projections. Therefore, we repeat the slicing process for many different slices and average the resulting distances. More formally, we compute the expected distance in one dimension between the projections of the respective distributions onto (uniformly) random directions on the unit sphere. As long as the distance of choice is a valid metric in one dimension, the sliced distance defined in this way is guaranteed to be a valid metric as well (Nadjahi et al., 2020, Proposition 1 (iii)). The most popular example of a sliced distance metric is the Sliced-Wasserstein (SW) distance (Fig. 2). However, we note that slicing has also been done for other distance measures, such as MMD with a specific choice of kernel (Hertrich et al., 2023) and mutual information (Goldfeld & Greenewald, 2021).
Figure 3: Computing Wasserstein distance. Two transport maps mapping the samples from a distribution $p_{1}$ (black) to samples from another distribution $p_{2}$ (blue), shown by arrows. The color of the arrow corresponds to the cost (Euclidean distance) between $x_{i}$ and $y_{i}$. (a) Randomly chosen transport map. (b) The optimal transport map, giving the smallest total cost. The total cost for the optimal map in (b) is the Wasserstein distance between these two sets of samples.
Wasserstein or earth mover distance
Intuitively, if two given probability distributions are thought of as two piles of dirt, the Wasserstein distance measures the (minimal) cost of “transporting” one pile of dirt to another (Panaretos & Zemel, 2019). The formal definition of the Wasserstein distance for continuous distributions, derived from optimal transport, is described in Section A.3. Here, we provide a more intuitive definition given a fixed set of samples from two distributions. Suppose we have samples $\{x_{1},\ldots,x_{N}\}\subset\mathbb{R}^{d}$ sampled from a distribution $p_{1}$ and $\{y_{1},...,y_{N}\}\subset\mathbb{R}^{d}$ sampled from another distribution $p_{2}$. Given any distance metric between two vectors in $\mathbb{R}^{d}$, $D(\cdot,\cdot)$, we can construct the cost matrix $C$, as the matrix of pairwise distances between the samples $x_{i}$ and $y_{j}$:
$C=\begin{bmatrix}D(x_{1},y_{1})&\dots&D(x_{1},y_{N})\\ \vdots&\ddots&\vdots\\ D(x_{N},y_{1})&\dots&D(x_{N},y_{N})\end{bmatrix}$ (1)
Recalling the earth-mover distance analogy, we want to map each $x_{i}$ to exactly one $y_{j}$, in such a way that the cost of doing so is minimized. The minimum transport map is then defining the Wasserstein distance (for the metric $D$) between the two empirical distributions. Throughout this work, we use the commonly used Euclidean metric, $L^{2}$, leading to the Wasserstein-2 and Sliced Wasserstein-2 distances. More precisely, we define a “transport map” to be a permutation matrix, $\pi\in\{0,1\}^{N\times N}$, which is a matrix with exactly one nonzero entry in each row. The entry $\pi_{ij}=1$ means that we transport the point $x_{i}$ to the point $y_{j}$. Then finding the transport map that minimizes the overall cost can be stated as
$\pi^{*}=\min_{\pi}\sum_{ij}\pi_{ij}C_{ij}.$ (2)
A randomly chosen transport map for small datasets in $\mathbb{R}^{2}$ is shown in Fig. 3a. Fortunately, the optimal solution to Eq. (2) can be solved exactly using the Hungarian method (Kuhn, 1955), leading to the assignment shown in Fig. 3b.
Slicing Wasserstein brings efficiency
Solving the optimal transport problem (Eq. (2)) with the Hungarian method has a time complexity of $O(N^{3})$ in the number of samples $N$ (although faster approximations exists, see Peyré et al. 2017). However, in the special case where the data is one-dimensional, the Wasserstein distance can be calculated by sorting the two datasets, obtaining the order statistics $\{x_{(1)},..,x_{(N)}\}$ and $\{y_{(1)},...,y_{(N)}\}$ and computing the sum of the distances $\sum_{i}D(x_{(i)},y_{(i)})$. This has a time complexity of $O(N\log(N))$. Thus, slicing the Wasserstein distance with one-dimensional projections becomes very efficient. While the value of the SW distance does not converge to the Wasserstein distance, even in the case of an infinite number of data samples and slices, the SW distance is a metric (in the mathematical sense) as long as $D$ is a metric on $\mathbb{R}^{d}$ and it acts as a lower bound to the Wasserstein distance (Nadjahi, 2021).
The Wasserstein distance and its sliced variant have several attractive properties: they can be computed differentiably; their computations do not rely heavily on choices of hyperparameters; and the sliced variant is very fast to compute. However, a disadvantage of the Wasserstein distance is its transparency: The numerical value of the Wasserstein distance has no intuitive interpretation due to its definition in terms of optimal transport maps. Additionally, per definition, the sliced variant is insensitive to differences within orthogonal subspaces of the slices. To still capture differences in all dimensions, naively, one would have to increase the number of slices (in the worst case) exponentially with the dimension, diminishing computational efficiency. However, other approaches exist to reduce this problem for nonlinear (Kolouri et al., 2019) or other specific (Deshpande et al., 2019; 2018) slices. Furthermore, slicing may also be relaxed to other kinds of data specific projections, such as Fourier features for stationary time series or locally connected projections for images (Du et al., 2023, Cazelles et al., 2020).
2.2 Classifier-based: Classifier Two-Sample Test (C2ST)
Figure 4: Classifier Two-Sample Test (C2ST). (a) The C2ST classifier problem: identifying the source distribution of a given sample. The optimal classifier predicts the higher-density distribution at every observed sample value, resulting in a majority of samples being correctly classified. (b) When probability densities of the distributions are not known, the optimal classifier is approximated by training a classifier, e.g., a neural network, to discriminate samples from the two distributions. (c) C2ST values vary from $0.5$ when distributions exactly overlap (left) to $1.0$ when distributions are completely separable (right).
The Classifier Two-Sample Test (C2ST) uses a classifier that discriminates between samples from two distributions (Fig. 4a) (Lopez-Paz & Oquab, 2016, Friedman, 2003). The distance between the distributions can then be quantified with various measures of classifier performance. The classification accuracy provides a particularly intuitive and interpretable measure of the similarity of the distributions. If the classification accuracy is 0.5, i.e., the classifier is at chance level, the distributions are indistinguishable to the classifier (Fig. 4c, left), while higher accuracy indicate differences in the distributions (Fig. 4c, middle). If the C2ST is 1.0, the two distributions have no (or very little) overlap in their supports (Fig. 4c, right). Given two distributions, the C2ST has a ‘true’ (optimal) value, which is the maximum classification accuracy attainable by any classifier (Fig. 4a). This optimal value can be computed if both distributions allow evaluating their densities, but this is not usually possible if only data samples are available. In that case, one aims to train a classifier, such as a neural network (Fig. 4b), that is as close to the optimal classifier as possible.
One of the main benefits of the C2ST is that its value is highly interpretable (the accuracy of the classifier). C2ST can also be used to test the statistical significance of the difference between two sets of samples. Unlike other measures, however, C2ST can be expensive to compute because it requires training a classifier and using the classification accuracy as a differentiable training objective is not straightforward (see Section A.5.1). Furthermore, the value is dependent on the capacity of the classifier, and hence on many hyperparameters such as classifier architecture or training procedure. In our experiments, we used a scikit-learn Multi-Layer-Perceptron classifier, combined with a five-fold cross-validation routine to estimate the accuracy returned (Pedregosa et al., 2011).
Common failure modes
As mentioned above, for any realistic scenario, the C2ST is computed by training a classifier. The resulting C2ST will only be a good measure of distance between real and generated data if the classifier is close or equal to the optimal classifier. To demonstrate the behavior of the C2ST if this is not the case, we fitted a Gaussian distribution to data that was sampled from a Mixture of Gaussians (Fig. 5a, left). The optimal C2ST between these two distributions is 0.65 (which can be computed because Gaussians and Mixtures of Gaussians allow evaluating densities). If the C2ST is estimated with a neural network, however, we observe that this C2ST can be systematically underestimated: for example, when only few samples from the data and generative model are available, the neural network predicts a C2ST of 0.5 (Fig. 5a, top right) – in other words, it predicts that the generative model and the data follow the same distribution. Similarly, if the neural network is not expressive enough, e.g. with too few hidden units, the classifier will return a low C2ST, around 0.5 (Fig. 5a, bottom right). These issues can make the C2ST easy to misuse: In many cases, reporting a low C2ST is desirable for generative models since it indicates that the model perfectly matches data, but one can achieve a low C2ST simply by not investing sufficient time into obtaining a strong classifier.
C2ST can remain very high even for seemingly good generative models
We previously argued that the C2ST is an interpretable measure – while this is generally true, the C2ST can sometimes be surprisingly high even if the generative model seems well aligned with the data. For example, when the generative model aligns very well with the data for every marginal, the C2ST can still be high if the data is high-dimensional (Fig. 5b). Because of this, it can be difficult to achieve low C2ST values on high-dimensional data. To further demonstrate this, we fitted a Gaussian distribution and a mixture of 20 Gaussian distributions to the ‘ones’ of the MNIST dataset. Although the Mixture of Gaussians (Fig. 5c, bottom row) looks better than a single Gaussian (Fig. 5c, middle row), both densities have a C2ST of 1.0 to the data (obtained with a ResNet on $\approx$4k held-out test datapoints).
Figure 5: Failure modes and behavior of C2ST. (a) Data (top left) and Gaussian maximum-likelihood estimate (bottom left). C2ST wrongly returns 0.5 (no difference between the densities) if too few samples are used (top right) or the neural network is poorly chosen (bottom right). (b) For high-dimensional densities, despite the marginals between data (black) and model (gray) seeming well-aligned, small differences (here a mean shift of 0.25 std. in every dimension) allow the classifier to more easily distinguish the distributions as dimensionality increases, yielding correct but surprisingly high C2ST. (c) On MNIST, the C2ST between data (top) and a Gaussian generative model (middle) as well as of a Mixture of Gaussians (MoG, bottom) is 1.0, although the MoG is perceptually more aligned with the data.
Other C2ST variants
While we focus on a standard C2ST definition by using classification accuracy as the C2ST distance (Lopez-Paz & Oquab, 2016), any other performance metric for binary classification could be used (Raschka, 2014). Kim et al. (2019) even argue that classic accuracy is sub-optimal due to the “binarization” of the class probabilities and proceeds to instead use the mean squared error between the predicted and ‘target’ value of 0.5. Other approaches instead construct a likelihood ratio statistic (Pandeva et al., 2022). Additionally, instead of using the estimated class probabilities, Cheng & Cloninger (2022) consider using the average difference in logits (i.e., activations in the last hidden layer). However, classification accuracy is still the most commonly used variant of C2ST.
2.3 Kernel-based: maximum mean discrepancy (MMD)
Figure 6: Maximum mean discrepancy (MMD). (a) Two example distributions $p_{1}(x)$, $p_{2}(x)$ and observed data $p_{true}(x)$ that we want to compare. (b) MMD can be defined as the difference between the expectations of some embedding function $\phi(x)$. If we take the identity as embedding ($\phi^{(1)}(x)=x$; left), we end up computing the differences between the means of the distributions, which are all equal for the three distributions. If we add a quadratic feature ($\phi(x)^{(2)}=[x,x^{2}]^{\mathsf{T}}$; right), we can distinguish distributions with different variances. Note that we still have $\mathsf{MMD}^{2}[\phi^{(2)},p_{2},p_{true}]=0$, despite $p_{2}$ being different from $p_{true}$ (c) Using the kernel trick we can avoid computing the embeddings all together but use implicit embeddings that capture all relevant features of the distributions.
MMD is a popular distance metric that is applicable to a variety of data domains, including high-dimensional continuous data spaces, strings of text as well as graphs (Borgwardt et al., 2006, Gretton et al., 2012a, Muandet et al., 2017). It has been used to evaluate generative models (Sutherland et al., 2021, Borji, 2019, Lueckmann et al., 2021) and also has the ability to indicate where the model and the true distribution differ (Lloyd & Ghahramani, 2015). The distance provided by MMD can straightforwardly be used to test whether the difference between two sets of high-dimensional samples is statistically significant (Gretton et al., 2012a).
To assess whether two set of samples are drawn from the same distribution, MMD makes use of a kernel function to (implicitly) embed the samples via an embedding function $\phi$, also called feature map. If we choose the right kernel, we can end up embedding our samples in a space where properties of the underlying distributions are easily compared. We will motivate the use of the kernel in MMD by illustrating different explicit embeddings before introducing the implicit embedding via a kernel $k$. Note that this explanation is inspired by Sutherland (2019).
In a first step, we can define MMD as the difference between the means of the embedding of two distributions $p_{1}$ and $p_{2}$:
$\mathsf{MMD}^{2}[\phi,p_{1},p_{2}]=\lVert\mathbb{E}_{p_{1}(x)}[\phi(x)]-\mathbb{E}_{p_{2}(y)}[\phi(y)]\rVert^{2},$
for any embedding function $\phi$.
If we want to compare samples of real numbers from two distributions $p_{1}$ and $p_{2}$ (Fig. 6a), we can think about different embedding functions $\phi$ to compare these. The simplest possible function $\phi^{(1)}:\mathbb{R}\rightarrow\mathbb{R}$ is the identity mapping $\phi^{(1)}(x)=x$ (Fig 6b, left). However, in this case the MMD will simply be the absolute difference between the means (first moments) of the distributions (for details, see Section A.2):
$\mathsf{MMD}[\phi^{(1)},p_{1},p_{2}]$ $=\lvert\mu_{p_{1}}-\mu_{p_{2}}\rvert.$
As both models and the true distribution in Fig. 6 have the same mean, this does not yet let us discriminate between them. If we now expand our embedding with a quadratic term, $\phi^{(2)}:\mathbb{R}\rightarrow\mathbb{R}^{2}$ as $\phi^{(2)}(x)=\begin{bmatrix}x\\ x^{2}\end{bmatrix}$ (Fig 6b, right), the MMD yields (for details, see Section A.2)
$\mathsf{MMD}^{2}[\phi^{(2)},p_{1},p_{2}]$ $=(\mu_{p_{1}}-\mu_{p_{2}})^{2}+(\mu_{p_{1}}^{2}+\sigma_{p_{1}}^{2}-\mu_{p_{2}}^{2}-\sigma_{p_{2}}^{2})^{2}.$
[…]
$\mathsf{MMD}^{2}[\phi,p_{1},p_{2}]=\mathbb{E}_{p_{1}(x),p_{1}^{\prime}(x^{\prime})}[\langle\phi(x),\phi(x^{\prime})\rangle]+\mathbb{E}_{p_{2}(y),p_{2}^{\prime}(y^{\prime})}[\langle\phi(y),\phi(y^{\prime})\rangle]-2\mathbb{E}_{p_{1}(x),p_{2}(y)}[\langle\phi(x),\phi(y)\rangle]$
We can now rewrite the inner product $\langle\phi(x),\phi(x^{\prime})\rangle$ in terms of a kernel function $k$: $\langle\phi(x),\phi(x^{\prime})\rangle=k(x,x^{\prime})$. Thus, if we can find a kernel for our feature map, we can avoid explicitly computing the features altogether but instead, we directly compute
$\mathsf{MMD}^{2}[k,p_{1},p_{2}]=\mathbb{E}_{p_{1}(x),p_{1}^{\prime}(x^{\prime})}[k(x,x^{\prime})]+\mathbb{E}_{p_{2}(y),p_{2}^{\prime}(y^{\prime})}[k(y,y^{\prime})]-2\mathbb{E}_{p_{1}(x),p_{2}(y)}[k(x,y)].$
Evaluating the kernel function instead of explicitly calculating the features is often called the kernel trick (Fig. 6c). If we can define a kernel whose corresponding embedding captures all, potentially infinitely many moments, we would have an MMD that is zero only if two distributions are exactly equal. These kernels are called characteristic (Section A.2,  Gretton et al. 2012a), and include the commonly used Gaussian kernel: $k_{G}(x,x^{\prime})=\exp(-\frac{\lVert x-x^{\prime}\rVert^{2}}{2\sigma^{2}})$ (see, e.g., Sriperumbudur et al. (2009) for other characteristic kernel choices).
MMD in practice
Typically the kernel version of MMD is used, which is straightforwardly estimated with its empirical (unbiased) estimate:
$\mathsf{MMD}^{2}=\frac{1}{m(m-1)}\sum_{i}^{m}\sum_{j\neq i}k(x_{i},x_{j})+\frac{1}{n(n-1)}\sum_{i}^{n}\sum_{j\neq i}k(y_{i},y_{j})-\frac{2}{mn}\sum_{i,j}k(x_{i},y_{j}).$
MMD in this form can be applied to many forms of data, as long as we can define a kernel, which can include graphs (Vishwanathan et al., 2010, Gärtner, 2003) or strings of text (Lodhi et al., 2002), in addition to vectors and matrices.
When we estimate the MMD with a finite number of samples, the selection of the right kernel and its parameters becomes crucial. For example, when using a Gaussian kernel, one has to choose the bandwidth $\sigma$. The MMD approaches zero if we take $\sigma$ to be close to zero (then $k_{G}(x,x^{\prime})=1\;\mathsf{if}\;x=x^{\prime}\;\mathsf{else}\;k_{G}(x,x^{\prime})\rightarrow 0$) or if $\sigma$ is large (then $k_{G}(x,x^{\prime})\rightarrow 1\;\forall\;x,x^{\prime}$(Gretton et al., 2012a). A common heuristic to remedy this parameter choice is picking the bandwidth based on the scale of the data. The median heuristic set the bandwidth to the median distance between points in the aggregate sample (Gretton et al., 2012a). Another common approach is based on cross-validation, or data splitting (Gretton et al., 2012a; b, Jitkrittum et al., 2016, Sutherland et al., 2021): The dataset is divided, with a hold-out set used for kernel selection, and the other part used for evaluating MMD. While the data splitting method does not involve any heuristic, it can lead to errors in MMD since it reduces the number of data points available for estimating the MMD. Recent work attempts to choose hyperparameters without employing data splitting or any heuristic (Biggs et al., 2023, Schrab et al., 2023, Kübler et al., 2022b; a).
While we aim in general for a kernel that captures the (dis)similarity between the data points well, such a kernel can be domain specific or specifically designed for downstream analysis tasks. The similarity between two strings (e.g., DNA sequences, text) can for instance be estimated by looking at the frequency of small subsequences (Leslie et al., 2001, Lodhi et al., 2002). It is furthermore possible to aggregate simpler kernels into a more expressive one (Gretton et al., 2012b), or to use a deep kernel (i.e., based on neural networks) that can exploit features of particular data modality such as images (Liu et al., 2020, Gao et al., 2021).
2.4 Network-based: Embedding-space measures
Distribution comparisons on structured data spaces, such as the set of natural images, present unique challenges. Such data is usually high-dimensional (high-resolution images) and contains localized correlations. Furthermore, images of different object classes (such as airplanes and dogs) share low-level features in the form of edges and textural details but differ in semantic meaning. Similar challenges occur for time-series data, natural language text, and other complex data type (Smith & Smith, 2020, Jeha et al., 2021).
Figure 7: Network-based metrics. Instead of directly computing distances in data space, complex data e.g., natural images of dogs sampled from $p_{1}(x)$ and aircraft sampled from $p_{2}(y)$, are jointly embedded into a vector space. The embedding function can, for example, be a deep neural network. The resulting distributions in feature space are then compared by a classical measure of choice $D$.
In this section, we take the example of natural images, but the presented framework generalizes to other data types. Naive distances would operate on a per-pixel basis, leading to scenarios where, for example, white dogs and black dogs are considered vastly different despite both being categorized as dogs. As we would like to have a distance measure that operates based on details relevant to the comparison we can leverage neural networks trained on a large image dataset that captures features ranging from low-level to high-level semantic details: While earlier layers in a convolutional neural network focus on edge detection, color comparison, and texture detection, later layers learn to detect high-level features, such as a dog’s nose or the wing of an airplane, which thought to be relevant for a meaningful comparison. Embedding-based distances use these activations of neural network layers as an embedding to compare the image distributions. The most popular distance in this class is the Fréchet Inception Distance (FID) (Heusel et al., 2017), used to evaluate generative models for images. The FID uses a convolutional neural net’s embeddings (specifically InceptionV3 (Szegedy et al., 2015a)) to extract the relevant features, applies a Gaussian approximation in the embedding space, and computes the Wasserstein distance on this approximation.
An FID-like measure, in essence, requires a suitable embedding network $f:\mathcal{X}\rightarrow\mathbb{R}^{d}$, where $f$ transforms the data from the original high-dimensional space $\mathcal{X}$ into a lower-dimensional, feature-rich representation in $\mathbb{R}^{d}$ (Fig. 7). Once the data samples are mapped into this reduced space through the embedding network, the two sets of embedded samples can be compared using the appropriate distance. When evaluating generative models for natural images, it is common to approximate the embedded distributions with Gaussian distributions by estimating their respective mean $\mu$ and covariances $\Sigma$. Under this Gaussian approximation, the squared Wasserstein distance (also known as the Fréchet distance) can be analytically computed as
$W^{2}((\mu_{1},\Sigma_{1}),(\mu_{2},\Sigma_{2}))=\|\mu_{1}-\mu_{2}\|^{2}+\text{Tr}\left(\Sigma_{1}+\Sigma_{2}-2\left(\Sigma_{1}\Sigma_{2}\right)^{\frac{1}{2}}\right).$
In principle, any appropriate metric can be used in place of the Fréchet distance. Jayasumana et al. (2023) show that the MMD (Section 2.3) can be better suited as a metric in the embedding space, as it perceptually matches human judgement on assessing image quality and coverage in generative models. A related and commonly used quality measure for images is the Inception Score (Salimans et al., 2016). In contrast to the FID, this measure uses the average InceptionV3 predicted class probabilities and compares them with the true marginal class distribution. Note that while both this score and the FID can agree with traditional distances (e.g. certain divergences), they might evaluate models differently (Betzalel et al., 2022); see Barratt & Sharma (2018) for further limitations of the Inception Score.
Limitations
One of the biggest limitations is the requirement of a suitable embedding net. Newer and more robust networks, such as the image network of the CLIP (Radford et al., 2021) vision-language model, provide better and semantically more consistent embeddings (Betzalel et al., 2022) than the InceptionV3 network. However, as the embedding network is generally non-injective, identical distributions in the embedding space may not necessarily translate to identical distributions in the original space. Previous research has demonstrated the FID’s sensitivity to preprocessing such as image resizing and compression (Parmar et al., 2022). Additionally, FID estimates are biased for finite sample sizes, making comparisons unreliable due to dependency on the generative model. However, methods to obtain a more unbiased estimate have been proposed ($\mathsf{FID}_{\infty}$; Chong & Forsyth 2020, Betzalel et al. 2022).
3 Comparison and scalability
When evaluating (or training) generative models, it is important to understand that different statistical distances penalize different features of the generated samples. They might for instance weigh differently how important it is to have large sample variability versus how well the modes of the true distribution are captured (as can be illustrated by optimizing a mis-specified model using the different distances, see Fig. S1 and  Theis et al. 2016). Thus, using different statistical distances to evaluate the quality of generative models can lead to different conclusions.
Figure 8: Scalability of different statistical distances with sample size and dimensionality. (a,b) Comparison of sample sets with varying sample size (between 50 and 4k samples per set) of a ‘true’ distribution either with a second dataset of the same distribution or with a sample set from an approximated/shifted distribution. We show the mean and standard deviation over five runs of randomly sampled data. Note that the subscript for MMD distances (bottom) denotes the bandwidth of the Gaussian kernel used for a given dataset and we report the squared distance for MMD. (a) Distances for the 2d-MoG example shown in Fig. 1 compared to samples from a unimodal Gaussian approximation with the same mean and covariance. (b) Distances for a ten-dimensional standard normal distribution, for which the first dimension is shifted by one for the shifted dataset. (c) Distances based on 10k samples from a standard normal distribution with varying dimensions (between 5 and 1000). As in (b), the first dimension is shifted by one for the ‘shifted’ dataset. We show the mean and standard deviation over five runs of randomly sampled data. One MMD bandwidth was selected for all n-dimensional datasets.
These differences can be even more pronounced in applications where we only have a limited amount of data points, e.g., identifying rare cell types (Marouf et al., 2020), or where we have very high-dimensional data, e.g., neural population recordings in neuroscience (Stringer et al., 2019). In such cases, one needs to ensure that the distance measures can reliably distinguish different distributions for the given sample set size while remaining computationally tractable. In the following sections, we investigate the sensitivity of the three presented distances which do not rely on embeddings, when it comes to distinguishing data sets with varying numbers of samples (Section 3.1) and varying numbers of dimensions (Section 3.2). As the absolute values of the distance measures are often hard to interpret and different measures are on different scales, we examined the relative distances by comparing two or more models to the true data. We, therefore, applied the distances to compare samples from a ‘true’ distribution against itself (intra-dataset) and against samples from another distribution that is either an approximation or a shifted version of the true distribution (inter-dataset).
In our experiments, we focused on inter- and intra-dataset comparisons for the following three datasets: First, we compared the two-dimensional Mixture of Gaussians (“2d-MoG”) dataset introduced in Fig. 1 with samples from a unimodal Gaussian approximation with the same mean and covariance as the true data samples. Second, we compared samples from a ten-dimensional standard normal distribution with a shifted normal distribution for which the first dimension is shifted by one (“10-dim Gaussian”). And last, we compared a standard normal distribution with varying dimensionality to shifted distributions for which we respectively shifted the first dimension for the inter-dataset comparison (“n-dim Gaussian”). We used default parameters for the SW and C2ST measures while adjusting the bandwidth parameter for the MMD measure for each of the three comparisons. Finally, as FID and other network-based distances do require an embedding network, we investigate the scaling properties specific to FID on the ImageNet dataset (in Section 3.3).
3.1 Varying number of samples
We explored the robustness of the distances to low sample sizes on the 2d-MoG and the 10-dim Gaussian dataset. We found that for the 2d-MoG dataset, all measures failed to reflect the dissimilarity of the distributions at the lowest sample size of 50 samples (Fig. 8a). However, C2ST’s behavior differs from MMD and SW, with C2ST indicating that the distributions are similar (C2ST $\approx$ 0.5) for both intra- (true) and inter-dataset (approx.) comparisons while the other two distances indicate they are different (distance $\neq 0$). The malfunction of C2ST can be harder to detect in such cases, compared to the one of MMD and SW. While the latter is easily identified by the incorrect intra-dataset results, the malfunction of C2ST is hard to detect for unknown distributions. For all measures, computed values quickly stabilized by a sample size of 1000 and yielded the expected results of low intra-dataset differences and high inter-dataset differences.
For the 10-dim Gaussian dataset, we observed that all distances can identify samples from the same distribution as more similar than samples from different distributions (Fig. 8b); while all three distances struggle with low sample sizes (see also Supp. Fig. S3, given enough samples, they all become robust, with no measure being clearly superior to the others. For the 2d-MoG experiment, more samples are required to clearly detect the difference between the two distributions (Fig. 8a) as compared to the case where the mean in one dimension is shifted (Fig. 8b). Intuitively, the larger the differences in the distributions we want to compare, the fewer samples we need to detect these differences (see additional experiments Supp. Fig. S3).
3.2 Varying dimensionality
We further tested how the distances scale with the data dimension using $n$-dimensional standard normal distributions. In the first experiment, the shifted distribution differed only in the first dimension, which was mean-shifted by one. As the dimensionality increases, all distances reliably indicate no difference in intra-dataset comparisons , but C2ST is the only measure that consistently identifies the inter-dataset difference (Fig. 8c). Note that this dataset is different from Fig. 5b, where mean shifts were applied to every dimension. When we changed the structure of the data distribution (e.g., by changing the mean of all dimensions or their variances, see Supp. Fig. S4), we observed a similar picture with some particularities: While the SW distance with a fixed number of slices has difficulties if the disparity between the distribution is only in one dimension, its performance drastically improves for differences in all dimensions, which is expected from the random projections SW is performing. C2ST seems to robustly detect differences even in high dimensions in these modified datasets, though previous experiments showed that this measure can be oversensitive to small changes in high dimensions (Fig. 5b,c). Lastly, MMD is not robust across different dimensions for a Gaussian kernel with a fixed bandwidth. We could make the MMD robust across dimensions by using median heuristic (Section 2.3), where we would increase the chosen bandwidth such that it stays on the order of the euclidean distance between datapoints. Note that MMD can in general be highly sensitive to its hyperparameters, and an appropriate value depends not only on the dimensionality of the data (Supp. Fig. S5 and S6), but also on the structure of the distributions (Supp. Fig. S4).
3.3 FID-like distance comparison on ImageNet
Figure 9: Comparison of distances for ImageNet. (a) A comparison between the ImageNet test set and samples generated by an unconditional diffusion model, with varying sample sizes. (b) Distance evaluation on dog classes (D) versus non-dog classes (~D), highlighting differences in image representation between these two categories of real data. (c) Distances between sets of randomly selected images vs varying number of included classes of images (from 10 to 1,000) from the test set, using synthetic samples created by a conditional diffusion model.
To also explore scaling properties of FID-like distances, we used images from the ImageNet dataset and embedded them with the InceptionV3 network (Deng et al., 2009, Szegedy et al., 2015b), following the implementation of Heusel et al. (2017). In addition to the 100,000 images in the ImageNet test dataset (1000 classes, 100 images per class), we generated high-quality synthetic samples using a state-of-the-art diffusion model as described by Dockhorn et al. (2022). We first produced 50,000 samples with the base unconditional version of this model. Using a conditional generative model, we additionally generated 100,000 class-conditional images (i.e., 100 per class), exactly matching the class distribution of the test-set. All of these images were embedded using the pre-trained InceptionV3 network (Szegedy et al., 2015a), transforming the raw images into a 2048-dimensional feature space.
Calculating the FID involves computing the mean and covariance of the distributions in the embedding space and then calculating the squared Wasserstein distance analytically. However, we broadened our evaluation by applying additional distances to the distributions in the embedding space. While SW distance, C2ST, and FID effectively highlight the greater dissimilarity of synthetic samples to real images (i.e., the ImageNet test set) ( Fig. 9a) even for low sample sizes, the distinctiveness of the FID becomes only apparent when analyzing more than around 2000 samples. Estimating the full covariance matrix of the 2048-dim features with fewer samples leads to degeneracy and, thus, to numerical issues computing the square root. Common implementations hence generally recommend using more than 2048 samples (Heusel et al., 2018a; b). As also shown by Jayasumana et al. (2023), Betzalel et al. (2022), the Gaussian assumption in the FID is violated and can lead to problematic behavior. In contrast, the other distances reliably estimate a larger inter-dataset distance in regimes with few samples.
In our subsequent analysis, we aimed to determine the effectiveness of various distances in discriminating between images from different classes by comparing the distances between different levels in the WordNet hierarchy (Miller, 1995). To this end, we focused on comparing images of dogs (D) with those of non-dog (~D) images (Fig. 9b). All investigated distances, except FID, were successful in identifying images from different classes as being more distinct than images from the same class. The FID comparison between data with multiple classes (~D vs. ~D) is higher than across dog classes and other classes (D vs. ~D). This shows that comparing two image classes can be problematic with FID, which is usually used to compare two distributions over many image classes (i.e. over natural images).
To examine the effects only including a subset of classes from the dataset (i.e., modes of the distribution), we employed a test set and a conditional synthetic dataset, each comprising 1,000 classes with 100 samples per class. Our analysis involved comparing the complete test set against synthetic datasets that included only subsets of classes (Fig. 9c). For comparison purposes and as a control measure, we also conducted a scenario where, instead of selectively excluding classes, we randomly removed an equivalent number of images from the dataset. This approach revealed that limiting the dataset to a small number of classes compromised the performance across all evaluated distances, in contrast to the outcomes observed when randomly excluding a subset of images. To achieve performance comparable to that observed with random removals, it was necessary to include at least 800 classes in the comparison. As the InceptionV3 network is, in essence, trained to classify ImageNet images (Szegedy et al., 2015a) (under certain regularization schemes), hence the extracted high-level features may also be very sensitive to class-dependent image features and not necessarily for general image quality. This behavior can be observed in Fig. 9c and was recently explored by Kynkäänniemi et al. (2022). By replacing InceptionV3 with other embedding networks (e.g. CLIP which is trained to match images to captions) this class sensitivity can be reduced (Kynkäänniemi et al., 2022).
3.4 Computational complexity
While we only considered the sensitivity and specificity of the different distances in the previous paragraphs, we want to highlight that they differ also in their computational complexity. With respect to number of samples $N$, MMD and FID have a complexity of $O(N^{2})$ (note that for MMD the computational cost can be reduced, potentially at the cost of making approximation; Gretton et al. 2012a, Zhao & Meng 2015, Cheng & Xie 2021, Bodenham & Kawahara 2023, Bharti et al. 2023, Gretton et al. 2012b). While SW distance scales with $O(N\log N)$ (Nadjahi, 2021), it is difficult to make principled assessments of the computational complexity for C2ST, as it is highly dependent on the chosen classifier. But as more samples lead to larger training and test datasets, sample size is likely to influence the compute time. Similarly, the computational complexity of computing these distances increases as the dimensionality of the data increases, with non-trivial scaling depending on the task and hyperparameters chosen.
Note that despite their differences, all presented distances are reasonably tractable in the settings of our experiments, whereas the scaling experiments might be computationally unfeasible for other distances or datasets. We therefore strongly recommend carefully considering the complexity of the measure before conducting experiments on high-dimensional or very large datasets.
4 Scientific applications
To demonstrate how the presented distances apply to evaluating generative models of scientific applications, we focus here on two examples: decision modeling in cognitive neuroscience and medical imaging. For each application, we used two generative models or simulators to sample synthetic data. We then compared the synthetic samples to real data using the discussed distances. To obtain baseline values for each distance, we computed distances between subsets of real data. For SW distance, MMD, and FID we anticipated values proximal to zero, while for the C2ST, we expected a value around 0.5. These baseline assessments provide a lower threshold of model fidelity to which we compared the deviation of model-generated samples.
4.1 Models of primate decision making
Figure 10: Comparing models of primate decision making. (a) Schematic of a drift-diffusion model (DDM), a classical neuroscientific model of decision making behavior. Overall evidence drives the model toward one of two choices (drift), but sensory and environmental noise result in random fluctuations in evidence integration (diffusion). (b) Distributions of primate decision times from the real dataset (black), and two fitted models of varying complexity: DDM1 (gray) and DDM2 (gold). (c) SW distance, C2ST, MMD (bandwidth=0.5) between subsets of the three distributions. FID is not applicable in these comparisons, because the data are one-dimensional distributions. Scatter-points indicate comparisons between ten random subsets from each dataset. Thick horizontal bars indicate median values.
We explored the fidelity of two generative models in replicating primate decision times during a motion-discrimination task (Roitman & Shadlen, 2002). We evaluated two versions of a Drift-Diffusion Model (DDM; Fig. 10a) (Ratcliff, 1978), a frequently used model in cognitive neuroscience. The two versions differ with respect to the drift rate, which is the speed and direction at which evidence accumulates towards a decision, and the decision boundaries, which determine how much evidence is needed to make a decision. Specifically the first version (DDM1) uses a drift rate that varies linearly with position and time, and decision boundaries that decay exponentially over time whereas the second version (DDM2) uses a drift rate and decision boundaries that are constant over time (for details, see Section A.9). We fitted each model against empirical primate decision times with the use of the pyDDM toolbox (Shinn et al., 2020), generated one-dimensional synthetic datasets, and compared each dataset to the actual primate decision time distributions. While the resulting distributions of decision times are visually similar (Fig. 10b), the DDM-generated distributions DDM1 and DDM2 are noticeably broader compared to the more tightly clustered real decision times. Moreover, the DDM1 distribution appears more similar to the real distribution than that of DDM2, which is shifted towards the left. As expected, the DDM1 model more precisely mimics the real data distribution, as compared to the DDM2, across the median values of the SW distance, MMD, and C2ST distances (Fig. 10c). For C2ST DDM1 and the real data distribution are even indistinguishable, with median C2ST values around $0.5$. This suggests that SW and MMD provide a more nuanced differentiation between the models.
4.2 Chest X-ray image generation
Figure 11: Comparing generated and real X-ray images. (a) Sketch of the embedded distributions of X-ray images from the three different datasets: real dataset (black), Progressive Growing Generative Adversarial Network (PGGAN) (grey), and Stable Diffusion (SD) model (gold). (b) Examples from real and generated X-ray images. Three full-view examples from each distribution and four examples magnifying the top right corner. (c) SW distance, C2ST, MMD (bandwidth=50), and FID between samples of the three distributions of embedded X-ray images. Scatter-points indicate comparisons between ten random subsets from each dataset. Thick horizontal bars indicate median values.
In the second application we turned to a high dimensional example, in which we compared synthetic X-ray images generated by a Progressive Growing Generative Adversarial Network (PGGAN) model  (Segal et al., 2021) and by a StableDiffusion (SD) model (Malik & Humair, 2023) to real chest X-ray images from the ChestX-ray14 dataset (Wang et al., 2017). Each image has a total dimension of $1024\times 1024$ pixels.
From visual inspection, we note two observations: First, the images produced by the SD model are clearer and sharper than either the real images or those generated by the PGGAN. Second, generated images contain unrealistic artifacts that distinguish them from real X-ray images (Fig. 11b). For instance, in real images, the top often contains annotations including e.g. patient id, side of the body, or the date the X-ray was taken. These textual elements often contain artifacts or, in case of SD, are completely unrealistic. To compare these high-dimensional images, we embedded them in a 512-dimensional embedding space using the CheXzero network (Tiu et al., 2022), a CLIP (Radford et al., 2021) network fine-tuned for chest X-ray images. We opted for using this specialized network instead of the standard InceptionV3 network as it might overcome biases introduced by classification task training (Kynkäänniemi et al., 2022). As expected, samples generated by PGGAN are closer to the real data across all distances compared to SD-generated data (Fig. 11c), likely due to the unrealistic sharpness and more obvious textual artifacts of the SD-generated images. However, C2ST is even high between PGGAN outputs and the real data, suggesting that the high-dimensionality of the data increases the sensitivity of this measure. Taken together, our results suggest that PGGAN is more accurate in generating realistic X-ray images compared to SD.
Our findings highlight that using different metrics can support different conclusions. For instance, C2ST suggests equality between DDM1 and real decision time data, whereas SW distance and MMD metrics indicate a larger difference between DDM1 and the real data. Similarly, in analyzing X-ray image generation, SW distance, MMD, and FID metrics suggest a high similarity between PGGAN-generated and real images, whereas C2ST indicates a strong difference. Thus, we want to highlight the importance of using multiple complementary distances for best results and understanding of model limitations.
5 Discussion
This work describes and explores four commonly applied sample-based distances representing different methodologies for defining statistical distance: Using low-dimensional projections (SW), obtaining a distance using classifiers (C2ST), using embeddings through kernels (MMD) or neural networks (FID). Despite their operational differences, they are all based on a fundamental concept: simplifying complex distributions into more manageable feature representations to facilitate comparison. Sliced distances effectively reduce multidimensional distributions to a set of one-dimensional distributions, where classical metrics are more easily applied or calculated. MMD uses kernels to (implicitly) project samples into a higher dimensional feature space, in which comparing mean values becomes more expressive. Classifier-based methods (C2ST) transform the task of distribution comparison into a classification problem; comparison is made by investigating how well a classifier can distinguish the distributions. Lastly, network-based distances, such as FID, explicitly map samples into a representative feature space and compare distributions directly within this space.
In the paragraphs below, we highlight the features and limitations of these investigated distances. Additionally, we discuss the relationships between these metrics and connect them to current related work.
Sliced Distances
Sliced distances stand out for their computational efficiency in evaluating distributional discrepancies. However, when distributions differ primarily in lower-dimensional subspaces, sliced distances might not detect these subtle differences without a large number of slices (see Fig. 8c). There are approaches to reduce this effect by considering other projections than simple linear slices, as described in Section 2.1. Due to its computational efficiency and differentiability, the SW distance can also be used as a loss function to train generative models (Wu et al., 2019, Deshpande et al., 2018; 2019, Liutkus et al., 2019, Vetter et al., 2024). In our experiments, the metric did show convincing results and in contrast to the MMD, C2ST, and FID, SW distance does not require to choose specific hyperparameters for which results can differ drastically. Although currently not extensively used in literature for evaluation, this makes the SW distance efficient, scalable, and an objective baseline for general distribution comparisons. Yet, this also makes it less flexible to adapt to specific features of interest. Although, the majority of research on sliced distances focus on sliced Wasserstein metrics, slicing other metrics is also possible. For a certain subset of choices, equivalence to MMDs can be established (Kolouri et al., 2019).
Classifier Two-Sample Test (C2ST)
C2ST distinguishes itself by producing an interpretable value: classification accuracy. This characteristic makes C2ST particularly appealing for practical applications, as it is easy to explain and interpret. A notable drawback is the computational demand associated with training a classifier, which can be substantial. Moreover, C2ST’s effectiveness is critically dependent on the selection and training of a suitable classifier. Interpreting results reported for C2ST requires knowledge of the classifier used and its appropriateness for the data at hand. Furthermore, automated training pipelines may encounter failures, such as when the trained classifier performs worse than chance, often due to overfitting to cross-validation folds (see also Section A.4). On the other hand, it is able to even detect subtle differences within two distributions in high dimensions. Even if there is a difference in only a single out of a thousand dimensions (for which SW distance and MMD might struggle), C2ST is able detect it (see Fig. 8c). This might be desirable, but can also be problematic. When comparing images, slight variations in a few pixels may not be visually noticeable, potentially making them unimportant to the researcher. In high-dimensional complex data, such slight variations are quite likely. Thus C2ST can be close to 1.0 in the high-dimensional setting, making it practically useless for evaluation (see Fig. 5c, 11c). The C2ST can be shown to be a MMD with a specific kernel function parameterized by the classifier (Liu et al., 2020).
MMD
The Maximum Mean Discrepancy is a strong tool for comparing two groups of data by looking at their average values in a special feature space. The effectiveness of MMD largely depends on the kernel function chosen (implicitly representing the feature space), which affects how well it can spot differences between various types of data. Inappropriate kernel choice can leave the metric insensitive to subtle differences in the distribution (Gretton et al., 2012b, Sriperumbudur et al., 2009) (see Fig. 8). The MMD can be estimated efficiently and is differentiable, and thus often used as a loss function for training generative models (Arbel et al., 2019, Li et al., 2017, Bińkowski et al., 2021, Briol et al., 2019). Yet, a kernel must satisfy certain criteria, e.g., positive definiteness, making the design of new kernel functions challenging. Such constraints are relaxed for FID-like metrics, which focus on explicit representations of the embedding, whereas (kernel) MMD instead focuses on implicit representations. One advantage, however, is that the implicit embedding allows for infinite dimensional feature spaces (through characteristic kernel functions). These can be proven to be able to discriminate any two distinct distributions, something that is impossible through explicit representations used by the FID. Recently,  Kübler et al. (2022a) proposed a method to estimate MMD via a witness function that determines MMD (Appendix A.2). This method is closely related to C2ST in that both two estimate a discrepancy among distributions via a classifier (Kübler et al., 2022a, Section 5).
Network-based
Network-based approaches for evaluating distributions focus on the analysis of complex data, emphasizing the importance of capturing high-level, semantically meaningful features. These methods leverage neural networks to project data into a lower-dimensional, feature-rich space where traditional statistical distances can be applied more effectively. This is particularly important for tasks where the visual or semantic quality of the data is important, making them a popular choice for assessing generative models in domains such as image and text generation. The primary challenge lies in the design of suitable network architectures that can extract relevant features for accurate distribution comparison. Even more important than for the C2ST, this network must be well-established and shared which is a necessary but not sufficient criterion (Chong & Forsyth, 2020) to compare different results. While such well-established defaults exist for images (Szegedy et al., 2015a, Radford et al., 2021), this is not the case for other domains. For example, the time series generation community did not yet establish a default, and embedding nets are either trained or chosen by the authors (Smith & Smith, 2020, Jeha et al., 2021). We showed that the class-sensitivity of FID (Section 3.3) tends to model collapse (such as GANs), but might not necessarily reflect general image quality. In fact, Betzalel et al. (2022), Kynkäänniemi et al. (2022), Jayasumana et al. (2023) found that relevant features sometimes can disagree with human judgment and that CLIP embeddings align more closely to what humans perceive as favorable or unfavorable. Yet, FID features have been shown to align much better with human perception than traditional metrics (Zhang et al., 2018).
Closing remarks
Ultimately, the choice of distance hinges on the nature of the data under consideration and the specific characteristics of it one aims to compare. Given a specific dataset and problem, one will likely have to look beyond the distances discussed in this paper. For example, in the realm of human-centric data like images and audio, the perceptual indistinguishability of distributions is important (Gerhard et al., 2013, Zhang et al., 2018). Time series data, characterized by its temporal structure, demand metrics that account for temporal shifts and variations in a manner that does not disproportionately penalize minor discrepancies in timing, such as Dynamic Time Warping (Müller, 2007), or by using frequency information (Hess et al., 2023). In general however, irrespective of the specific use-case, it is advisable to use multiple distances in order to obtain a full picture, as using a single distances individually could support competing conclusions about the model that is to be evaluated.
Throughout this paper we have explained and analyzed four approaches of measuring statistical distance. While this is only a small subset of all possible distances available, we hope to have provided the foundational knowledge with which researchers can find, understand and interpret statistical distances specific to their own scientific application.
Code availability
All code for replicating and running our analysis is available at: https://github.com/mackelab/labproject. Note that for implementations of the C2ST we largely made use of those from the simulation-based-inference package (Tejero-Cantero et al., 2020).
Author Contributions
Conceptualization: AD, AS, CS, GM, JKL, LH, MD, MG, MP, SB, ZS
Software: AD, AES, AS, FG, FP, GM, JK, JKL, JV, LH, MD, MG, MP, RG, SB, ZS
Writing — Original Draft: AS, CS, FG, FP, GM, JK, JV, LH, MD, MG, MP, RG, RR, SB, ZS
Writing — Review & Editing: AD, AS, CS, FP, GM, JHM, JKL, JV, LH, LU, MG, MP, SB, ST, ZS
Visualization: AD, AES, AS, FP, GM, JKL, JV, MD, MG, MP, RG, RR, ZS
Project administration: AS, MG, MP, SB
Funding acquisition: JHM
Acknowledgments
We thank Eiki Shimizu and Nastya Krouglova for reviewing an early version of this paper. This project was inspired by a similarly structured project by the research group led by Karim Jerbi.
This work was supported by the German Research Foundation (DFG) through Germany’s Excellence Strategy (EXC-Number 2064/1, PN 390727645) and SFB1233 (PN 276693517), SFB 1089 (PN 227953431), SPP2041 (PN 34721065), the German Federal Ministry of Education and Research (Tübingen AI Center, FKZ: 01IS18039; DeepHumanVision, FKZ: 031L0197B; Simalesam: FKZ 01|S21055A-B), the Carl Zeiss Foundation (Certification and Foundations of Safe Machine Learning Systems in Healthcare), the Else Kröner Fresenius Stiftung (Project ClinbrAIn), and the European Union (ERC, DeepCoMechTome, 101089288). ST was partially supported by Grant-in-Aid for Research Activity Start-up 23K19966 and JST CREST JPMJCR2015. The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting SB, MD, MG, JK, JL, GM, MP, RR, AS, ZS and JV.
Appendix A Appendix
A.1 Generative Models in Science
Table S1: Example generative models in science. Simulators include mechanistic models, DNN based models e.g. VAEs, GANs, Diffusion models. This is not an exhaustive list regarding disciplines using generative models nor generative models used in the listed disciplines.
ML model Simulator
natural science biology
[…]
geoscience
[…]
chemistry
[…]
astronomy
[…]
A.2 Details about Maximum Mean Discrepancy
Here we provide different formulations and examples of MMD.
Definition A.1 (Feature Map Definition of MMD)
$\mathsf{MMD}^{2}[\phi,p_{1},p_{2}]=\lVert\mathbb{E}_{p_{1}(x)}[\phi(x)]-\mathbb{E}_{p_{2}(y)}[\phi(y)]\rVert_{\mathcal{H}}^{2},$ (3)
where $p_{1}(x)$ and $p_{2}(y)$ are the probability distributions of random variables $x,y\in\mathcal{X}$, and $\phi:\mathcal{X}\rightarrow\mathcal{H}$.
Generally $\mathcal{X}$ and $\mathcal{H}$ are defined as a topological space, and the reproducing kernel Hilbert space (RKHS), respectively, but readers can simply think of the euclidean space $\mathbb{R}^{N}$ for the first examples in the main text.
For the identity feature map $\phi^{(1)}:\mathbb{R}\rightarrow\mathbb{R},\phi^{(1)}(x)=x$, MMD can be computed as
$\mathsf{MMD}^{2}[\phi^{(1)},p_{1},p_{2}]$ $=\lVert\mathbb{E}_{p_{1}(x)}[x]-\mathbb{E}_{p_{2}(y)}[y]\rVert_{\mathbb{R}}^{2}$
$=(\mathbb{E}_{p_{1}(x)}[x]-\mathbb{E}_{p_{2}(y)}[y])^{2}$
$\mathsf{MMD}[\phi^{(1)},p_{1},p_{2}]$ $=\lvert\mu_{p_{1}}-\mu_{p_{2}}\rvert.$
And for the quadratic polynomial feature map $\phi^{(2)}:\mathbb{R}\rightarrow\mathbb{R}^{2},\phi^{(2)}(x)=\begin{bmatrix}x\\ x^{2}\end{bmatrix}$, MMD can be computed as
$\mathsf{MMD}^{2}[\phi^{(2)},p_{1},p_{2}]$ $=\lVert\mathbb{E}_{p_{1}(x)}[\begin{bmatrix}x\\ x^{2}\end{bmatrix}]-\mathbb{E}_{p_{2}(y)}[\begin{bmatrix}y\\ y^{2}\end{bmatrix}]\rVert_{\mathbb{R}}^{2}$
$=\lVert\begin{bmatrix}\mu_{p_{1}}\\ \mu_{p_{1}}^{2}+\sigma_{p_{1}}^{2}\end{bmatrix}-\begin{bmatrix}\mu_{p_{2}}\\ \mu_{p_{2}}^{2}+\sigma_{p_{2}}^{2}\end{bmatrix}\rVert_{\mathbb{R}}^{2}$
$\mathsf{MMD}^{2}[\phi^{(2)},p_{1},p_{2}]$ $=(\mu_{p_{1}}-\mu_{p_{2}})^{2}+(\mu_{p_{1}}^{2}+\sigma_{p_{1}}^{2}-\mu_{p_{2}}^{2}-\sigma_{p_{2}}^{2})^{2}.$
Definition A.2 (Kernel Definition of MMD)
$\mathsf{MMD}^{2}[\phi,p_{1},p_{2}]$ $=\lVert\mathbb{E}_{p_{1}(x)}[\phi(x)]-\mathbb{E}_{p_{2}(y)}[\phi(y)]\rVert_{\mathcal{H}}^{2}$
$=\langle\mathbb{E}_{p_{1}(x)}[\phi(x)],\mathbb{E}_{p_{1}(x)}[\phi(x)]\rangle_{\mathcal{H}}+\langle\mathbb{E}_{p_{2}(y)}[\phi(y)],\mathbb{E}_{p_{2}(y)}[\phi(y)]\rangle_{\mathcal{H}}-2\langle\mathbb{E}_{p_{1}(x)}[\phi(x)],\mathbb{E}_{p_{2}(y)}[\phi(y)]\rangle_{\mathcal{H}}$
$=\mathbb{E}_{p_{1}(x),p_{1}^{\prime}(x^{\prime})}[\langle\phi(x),\phi(x^{\prime})\rangle_{\mathcal{H}}]+\mathbb{E}_{p_{2}(y),p_{2}^{\prime}(y^{\prime})}[\langle\phi(y),\phi(y^{\prime})\rangle_{\mathcal{H}}]-2\mathbb{E}_{p_{1}(x),p_{2}(y)}[\langle\phi(x),\phi(y)\rangle_{\mathcal{H}}]$
$\mathsf{MMD}^{2}[k,p_{1},p_{2}]$ $=\mathbb{E}_{p_{1}(x),p_{1}^{\prime}(x^{\prime})}[k(x,x^{\prime})]+\mathbb{E}_{p_{2}(y),p_{2}^{\prime}(y^{\prime})}[k(y,y^{\prime})]-2\mathbb{E}_{p_{1}(x),p_{2}(y)}[k(x,y)].$
The definition of MMD can be rewritten through the notion kernel mean embedding. For given distribution $p(x)$, the kernel mean embedding $\mathbb{E}_{p(x)}[k(x,u)]\in\mathcal{H}$ is an element in RKHS that satisfies $\langle\mathbb{E}_{p(x)}[k(x,u)],f(u)\rangle_{\mathcal{H}}=\mathbb{E}_{p(x)}[f(x)]$ for any $f\in\mathcal{H}$ with argument $u\in\mathcal{X}$. The embedding $\mathbb{E}_{p(x)}[k(x,u)]$ is known to be determined uniquely if a corresponding kernel is bounded, i.e. $\|k(x,x^{\prime})\|_{\mathcal{H}}<\infty$ for any $x$. Then, as shown in Gretton et al. (2012a), $\mathsf{MMD}^{2}$ can be represented as
$\mathsf{MMD}^{2}[k,p_{1},p_{2}]=\|\mathbb{E}_{p_{1}(x)}[k(x,u)]-\mathbb{E}_{p_{2}(y)}[k(y,u)]\|^{2}_{\mathcal{H}}.$
MMD can also be defined more generally as the integral probability metric.
Definition A.3 (Supremum Definition of MMD)
$\mathsf{MMD}[\mathcal{F},p_{1},p_{2}]=\sup_{f\in\mathcal{F}}(\mathbb{E}_{p_{1}(x)}[f(x)]-\mathbb{E}_{p_{2}(y)}[f(y)]).$ (4)
Here, $\mathcal{F}$ is a class of functions $f:\mathcal{X}\rightarrow\mathbb{R}$. Where we take $\mathcal{F}$ as the unit ball in an RKHS $\mathcal{H}$ with associated kernel $k(x,x^{\prime})$ (Gretton et al., 2012a), the function that attains supremum (the witness function) is
$f(u)=\frac{\mathbb{E}_{p_{1}(x)}[k(x,u)]-\mathbb{E}_{p_{2}(y)}[k(y,u)]}{\|\mathbb{E}_{p_{1}(x)}[k(x,u)]-\mathbb{E}_{p_{2}(y)}[k(y,u)]\|_{\mathcal{H}}}.$
Assigning $f(u)$ into (Eq. (4)), we have
$\mathsf{MMD}^{2}[\mathcal{F},p_{1},p_{2}]$ $=\Bigl{(}\sup_{f\in\mathcal{F}}(\mathbb{E}_{p_{1}(x^{\prime})}[f(x^{\prime})]-\mathbb{E}_{p_{2}(y^{\prime})}[f(y^{\prime})])\Bigr{)}^{2}$
$=\Bigl{(}\frac{\mathbb{E}_{p_{1}(x),p_{1}^{\prime}(x^{\prime})}[k(x,x^{\prime})]+\mathbb{E}_{p_{2}(y),p_{2}^{\prime}(y^{\prime})}[k(y,y^{\prime})]-2\mathbb{E}_{p_{1}(x),p_{2}(y)}[k(x,y)]}{\|\mathbb{E}_{p(x)}[k(x,u)]-\mathbb{E}_{p_{2}(y)}[k(y,u)]\|_{\mathcal{H}}}\Bigr{)}^{2}$
$=\Bigl{(}\frac{\|\mathbb{E}_{p_{1}(x)}[k(x,u)]-\mathbb{E}_{p_{2}(y)}[k(y,u)]\|_{\mathcal{H}}^{2}}{\|\mathbb{E}_{p_{1}(x)}[k(x,u)]-\mathbb{E}_{p_{2}(y)}[k(y,u)]\|_{\mathcal{H}}}\Bigr{)}^{2}$
$=\|\mathbb{E}_{p_{1}(x)}[k(x,u)]-\mathbb{E}_{p_{2}(y)}[k(y,u)]\|^{2}_{\mathcal{H}},$
which is equal to the kernel definition of MMD.
Definition A.4 (Characteristic Kernel)
A kernel is called characteristic when the kernel mean embedding
$p(x)\mapsto f(u)=\mathbb{E}_{p(x)}[k(x,u)]\in\mathcal{H}$
is injective (Sriperumbudur et al., 2011, Fukumizu et al., 2008).
This means that, if a characteristic kernel is used, the embedding into the RKHS can uniquely preserve all information about a distribution. In our evaluation, we utilize the Gaussian kernel, one of the well-known characteristic kernels. Another example of a characteristic kernel is the Laplacian kernel, which is defined by
$k(x,x^{\prime}):=\text{exp}\bigr{(}-\beta|x-x^{\prime}|\bigl{)}.$
Note that linear and polynomial kernels are not characteristic, while they are quite popular in natural language processing.
A.3 Details about Wasserstein and Sliced-Wasserstein distances
Formally, the Wasserstein distance is described in measure-theoretic terms. We first state this definition, and then provide an accessible interpretation in the common case that the measures have well-defined probability density functions.
Let $(M,\rho)$ be a Polish space, $\mu,\nu\in P(M)$ be two probability measures, and let $q\in[1,+\infty)$. Then the Wasserstein-$q$ distance between $\mu$ and $\nu$ is defined as
$W_{q}(\mu,\nu)=\left(\inf\limits_{\gamma\in\Gamma(\mu,\nu)}\mathbb{E}_{x_{1},x_{2}\sim\gamma}\rho(x,y)^{q}\right)^{1/q},$
where $\Gamma(\mu,\nu)$ is the set of all couplings between $\mu,\nu$.
Consider the case where $M\subseteq\mathbb{R}^{d}$, $\rho$ is the $q$-norm $||\cdot||_{q}$, and $\mu$ and $\nu$ have well-defined probability density functions $p_{1}$ and $p_{2}$ respectively. Then the Wasserstein-$q$ norm can be written as
$W_{q}(p_{1},p_{2})=\inf_{\gamma\sim\Gamma(p_{1},p_{2})}\left(\mathbb{E}_{x_{1},x_{2}\sim\gamma}\,||x_{1}-x_{2}||_{q}^{q}\right)^{\frac{1}{q}},$
where $\Gamma(p_{1},p_{2})$ is the set of all couplings, that is all possible “transportation plans”, between $p_{1}$, $p_{2}$. $\gamma\in\Gamma(p_{1},p_{2})$ is a joint distribution over $(x_{1},x_{2})$ with respective marginals $p_{1}$ and $p_{2}$ over $x_{1}$ and $x_{2}$.
The Sliced-Wasserstein distance is similarly defined in measure-theoretic terms for the measures $\mu$ and $\nu$. We refer the reader to Nadjahi (2021) for details. In the less general case described above, we can similarly provide a more intuitive definition. In particular, the random projection directions described in 2.1 are uniformly random vectors $u\in\mathbb{S}^{d-1}$, the unit sphere in $\mathbb{R}^{d}$. Projecting the distributions $p_{1}$ and $p_{2}$ onto $u$ induces one-dimensional distributions $p_{1}^{u}$ and $p_{2}^{u}$ with samples $u^{\mathsf{T}}x_{i}$, where $x_{1}\sim p_{1}$ and $x_{2}\sim p_{2}$. The Sliced-Wasserstein-$q$ distance can then be written as
$SW_{q}(p_{1},p_{2})=\mathbb{E}_{u\sim\mathcal{U}(\mathbb{S}^{d-1})}[W_{q}(p_{1}^{u},p_{2}^{u})],$
where $\mathcal{U}(\mathbb{S}^{d-1})$ is the uniform distribution over vectors on the unit sphere $\mathbb{S}^{d-1}$.
A.4 C2ST scores below 0.5
In practice the C2ST score can sometimes turn out to be below .5. That is, the trained classifier performs systematically worse than a random classifier. A potential reason for this effect is the existence of near duplicates or copies between the two different sets. Before training the classifier, these duplicates are assigned opposite class labels. When a given pair of such duplicates is then split into one that belongs to the training set and one that belongs to the test set, the classifier is biased towards predicting the wrong class for the duplicate in the test set. This effect is particularly noticeable if the classifier was not carefully regularized during training, and thus memorized the class label of the duplicate in the training set.
A.5 Sliced-Wasserstein, MMD and C2ST as optimization target
Figure S1: Trade-offs illustrated through optimization of a miss-specified model. We fitted a misspecified model (a two-dimensional Gaussian) by using different distances to a multi-modal distribution (similar to Theis et al. 2016). Note, that for a well-specified model each distance would give a perfect fit (Section A.5.2). In the optimization we minimized the SW distance, the C2ST classifiability, and the MMD with a Gaussian Kernel (details in Section A.5.1). Plotted are the contour lines of .25, .75, 1, and 2.5 standard deviations of the fitted Gaussians. The model optimised with SW is mass-covering: it covers both modes and therefore also assigns density to low-density regions of the true distribution, thus producing varied, but potentially unlikely samples. The models optimised with C2ST and MMD are mode-seeking: they have high densities only in the largest mode of the true distribution, and thus produces likely, but unvaried samples.
A.5.1 Fitting a Gaussian with gradient descent
We provide an illustrative example of which distributions are obtained when using Wasserstein, MMD and C2ST distances as a goodness of fit criterion, for both a miss-specified example, in Fig. S1 and well-specified example Fig. S2. We can see that in the miss-specified example different distances make different trade-offs, for example whether they are mode-seeking, and produce likely but unvaried samples, or are mode-covering, where they produce varied, but also potentially unlikely samples.
For Wasserstein, optimisation has been studied more formally in previous work (Bernton et al., 2019, Yi & Liu, 2023). Wasserstein was used as training objective in Arjovsky et al. (2017) and MMD was used as training objective in Bińkowski et al. (2021), Dziugaite et al. (2015), Li et al. (2015). Optimizing the C2ST classifier at the same time as the parameters of our generative model is similar to training a GAN (Goodfellow et al., 2014), but for simplicity we instead optimized for the closed-form optimal C2ST as for this toy example we have access to true densities. While FID can be used as optimization target in principle (Mathiasen & Hvilshøj, 2021), its applicability to our toy example here is less obvious, so we excluded it here.
In order to fit the (miss-specified) Gaussian model,
$p(x)=\mathcal{N}(\mu,\Sigma)$
to the ground truth distribution $p_{\text{true}}$, which is a mixture of Gaussians, we proceed as follows. Let $CC^{\mathsf{T}}$ denote the Cholesky decomposition of $\Sigma$. We compute gradients with respect to $\mu$ and $C$ by using the reparameterization trick; by generating samples as $\mu+C{\epsilon}$, with ${\epsilon}\sim\mathcal{N}(0,\mathsf{I})$.
For Wasserstein we used as loss the Sliced Wasserstein distance, for MMD, we used a Gaussian Kernel with bandwidth set according the median heuristic.
For C2ST, we can evaluate the probability densities of samples from both the learned Gaussian and the ground truth mixture of Gaussians, so we minimize the accuracy of the closed-form optimal classifier. For each sample, we evaluate the log-probability density of the sample under each distribution, softmax the two resulting values, and use those as the classifier predicted probabilities. We then use binary cross-entropy as the loss function.
We used the ADAM optimizer (Kingma & Ba, 2015), with learning rate=0.01 and default momentums, using 2500 epochs of 10000 samples.
A.5.2 Fitting a mixture of Gaussians with Expectation-Maximisation
We also include an example where the model we fit is well-specified, which in this case means it is also a mixture of two Gaussians (Fig. S2. As directly optimizing a mixture distribution with gradient descent is not straightforward, we used an Expectation-Maximization algorithm (where we use the distances instead of the log-likelihood in the maximisation step)
Our model is specified by
$p(x)=w_{1}\mathcal{N}(\mu_{1},\Sigma_{1})+w_{2}\mathcal{N}(\mu_{2},\Sigma_{2})$
which we can write as a latent-variable model, where the latent variables are the cluster assignments:
$p(x)=\sum_{k=1}^{2}p(x|z=k)p(z=k),$
with $p(x|z=k)=\mathcal{N}(\mu_{k},\Sigma_{k})$ and $p(z=k)=w_{k}$.
We then iteratively performed the following two steps to optimise the model.
E-step: For each of the $N$ datapoints $\hat{x}_{i}$ from $p_{\text{true}}$, we calculated the probability of it belonging mixture component 1 or 2:
$p(z=k|\hat{x}_{i})=\frac{p(\hat{x}_{i}|z=k)p(z=k)}{\sum_{k^{\prime}}^{2}p(\hat{x}_{i}|z=k^{\prime})p(z=k^{\prime})}.$
M-step: We updated the mixture weights according to:
$w_{k}=\frac{1}{N}\sum_{i}^{N}p(z=k|\hat{x}_{i})$
Next, for each of the $N$ datapoints, we first sampled a cluster assignment according to $z_{i}\sim p(z|\hat{x}_{i})$. Then for each group of $N_{k}$ datapoints assigned to cluster $k$ we sampled $N_{k}$ times according to $x_{i}\sim p(x|z_{i})$, again using the reparameterisation trick. As before we computed the loss using a statistical distance, now separately for the two groups of samples assigned to either mixture component, and used gradient to optimise $\mu_{k},\Sigma_{k}$
Again, we used the ADAM optimizer (Kingma & Ba, 2015), with learning rate=0.01 and default momentums, using 2000 epochs of 5000 samples.
Figure S2: Optimising distances in a well-specified model setting When fitting a well-specified model (here, a mixture of two Gaussians), by using different distances in the loss, we can see that each model converges to the global optimum. Plotted are the contour lines of .25, .75, 1, and 2.5 standard deviations of the fitted Gaussians multiplied by their corresponding mixture weight.
A.6 Additional scaling experiments with different sample size budgets and ranges
In Fig. 8, we evaluated the robustness of the measures against the number of samples and the dimensionality of the data. We observed notably poor performance of the measures in scenarios with limited data. Here, we further examine the performance of the distances across datasets of varying sample sizes, particularly for small sample set sizes, ranging from only 8 to 80 samples per set (Fig. S3). We examine three distinct data configurations where the distinction between the true and approximated distributions progressively increases from subpanels S3 a to c. Across all distances, it becomes evident that the larger the disparity between the two distributions, the fewer samples are needed for differentiation. In the experiment where all dimensions are mean-shifted by one, a sample size of 8 is sufficient to distinguish between the distributions. However, for less distinct distributions, such as the unimodal Gaussian or a mean-shift by one in only one dimension (Supp. Fig. S3 a, b), all distances exhibit poor performance in distinguishing between the distributions.
Figure S3: The larger the difference between two distributions the fewer samples suffice to tell the true and shifted distribution apart. We compare sample sets with varying sample sizes (between 8 and 80 samples per set) of a ’true’ distribution either with a second dataset of the same distribution or with a sample set from an approximated/shifted distribution. We show the mean and standard deviation over five runs of randomly sampled data. (a) Distances for the 2d-MoG example shown in Fig. 1 compared to samples from a unimodal Gaussian approximation with the same mean and covariance. (b) Distances for a ten-dimensional standard normal distribution, for which the first dimension is shifted by one for the shifted example. (c) Distances for a ten-dimensional standard normal distribution, for which all dimensions are shifted by one for the shifted example.
A.7 Additional scaling experiments for different dimensionality of the data
When comparing the robustness of the measures with respect to the dimensionality of the data in Fig. 8, we observed a degradation in the ability to distinguish between distributions as dimensionalities increased. Notably, only the C2ST measure retained the capability to distinguish between the two distributions in higher dimensions which is aligned with the intuition that a classifier can easily pick up on differences in a single dimension. Extending this analysis, Fig. S4 presents similar experiments conducted on datasets where we compare an n-dimensional standard normal distribution with one where either all dimensions are mean-shifted by one (thus aligning with the C2ST experiment in Fig. 5b) or where all variances are increased by one. Fig. S4a corresponds to the experiment outlined in Fig. 8c on dimensionality. The bandwidth parameter for the MMD distance has been adjusted to suit the particular data configuration and is represented by the integer in the y-axis label. Generally, we notice that the Sliced-Wasserstein distance and MMD face difficulties in higher-dimensional spaces, especially when handling distributions that are only slightly distinct if the respective hyperparameters are kept constant across dimensions. In contrast, the C2ST distance consistently demonstrates good performance across all three experiments and for all ranges of dimensions.
Figure S4: The impact of dimensionality varies across distances, with certain distances facing particular challenges in higher dimensional spaces. We compare sample sets of varying dimensionality (between 5 and 1000) of a ’true’ distribution either with a second dataset of the same distribution or with a sample set from an approximated/shifted distribution. The sample size is fixed to $10$k for all experiments and we show the mean and standard deviation over five runs of randomly sampled data. The bandwidth parameter in Gaussian Kernel MMD is set to 10 for all experiments. (a) Distances for a sample set from an n-dimensional standard normal distribution, for which the first dimension is shifted by one. (b) Distances for a sample set from an n-dimensional standard normal distribution, for which all dimensions are shifted by one. (c) Distances for a sample set from an n-dimensional standard normal distribution, for which variances are increased by one for all dimensions.
A.8 Sensitivity of the MMD bandwidth parameter
Figure S5: The bandwidth parameter in Gaussian Kernel MMD is a sensitive parameter that requires careful selection for each dataset. The sample size is fixed to $10$k for all experiments and we show the mean and standard deviation over five runs of randomly sampled data. (a) $MMD^{2}$ distance with varying bandwidth parameters between 0.1 and 5 for the 2d-MoG example compared to samples from a unimodal Gaussian approximation with the same mean and covariance. (b) $MMD^{2}$ distance with varying bandwidth parameters between 0.1 and 20 for a 10-dimensional standard normal distribution, for which the first dimension is shifted by one for the shifted example. (c) $MMD^{2}$ distance with varying bandwidth parameters between 1 and 40 for a 100-dimensional standard normal distribution, for which the first dimension is shifted by one for the shifted example.
Figure S6: Evaluating the effect of varying bandwidth parameters in Gaussian Kernel MMD for different sample sizes and dataset dimensionalities. (a,b) We compare sample sets with varying sample sizes (between 50 and 4k samples per set) of a ’true’ distribution either with a second dataset of the same distribution or with a sample set from an approximated/shifted distribution. We show the mean and standard deviation over five runs of randomly sampled data. (a) $MMD^{2}$ distance with varying bandwidth parameters (0.5, 1, 2) for the 2d-MoG example shown in Fig. 1 compared to samples from a unimodal Gaussian approximation with the same mean and covariance. (b) $MMD^{2}$ distance with varying bandwidth parameters (1, 3, 5) for a ten dimensional standard normal distribution, for which the first dimension is shifted by one for the shifted example. (c) $MMD^{2}$ distance with varying bandwidth parameters (5, 10, 40) based on 10k samples from a standard normal distributions with varying dimensions (between 5 and 1000). As in (b) the first dimension is shifted by one for the ’shifted’ dataset. Here we show one run due to computational costs.
A.9 Details about scientific application examples
For the motion discrimination task, we used the decision times of a single animal during both correct and erroneous trials with dot motion coherence of 12.8%, leading to a one-dimensional dataset of 587 samples. DDMs were implemented using the pyDDM toolbox (Shinn et al., 2020). DDM1 used a linear drift and exponential decision boundaries. In contrast, DDM2 used a constant drift and a constant decision boundary. Both were sampled 1,000 times to create the two synthetic datasets. The real chest X-ray dataset consists of 70,153 samples, the generated datasets from PGGAN and SD consist of 10,000 and 2,352 samples respectively.
In both applications we computed metrics between pairs of 10 random subsets from the compared distributions (scatter points on the violin plots). We computed the MMD with a bandwidth of 50 for the medical imaging datasets and a bandwidth of 0.5 for the decision time dataset.
[…]

### v2

@@
 How do we evaluate the samples these models generate?
-This work aims to provide an accessible entry point to understanding popular notions of statistical distances, requiring only foundational knowledge in mathematics and statistics.
+This work aims to provide an accessible entry point to understanding popular sample-based statistical distances, requiring only foundational knowledge in mathematics and statistics.
 We focus on four commonly used notions of statistical distances representing different methodologies: Using low-dimensional projections (Sliced-Wasserstein; SW), obtaining a distance using classifiers (Classifier Two-Sample Tests; C2ST), using embeddings through kernels (Maximum Mean Discrepancy; MMD), or neural networks (Fréchet Inception Distance; FID).
 We highlight the intuition behind each distance and explain their merits, scalability, complexity, and pitfalls.
-To demonstrate how these distances are used in practice, we evaluate generative models from different scientific domains, namely a model of decision making and a model generating medical images.
+To demonstrate how these distances are used in practice, we evaluate generative models from different scientific domains, namely a model of decision-making and a model generating medical images.
 We showcase that distinct distances can give different results on similar data.
@@
 Generative models that produce samples of complex, high-dimensional data, have recently come to the forefront of public awareness due to their utility in a variety of scientific, clinical, engineering, and commercial domains (Bond-Taylor et al., 2021).
-Prominent examples include StableDiffusion and DALL-E for generating photo-realistic images (Rombach et al., 2022), WaveNet (Oord et al., 2016) for audio synthesis, and Generative Pre-trained Transformer (GPT;  Radford et al.
+Prominent examples include StableDiffusion (SD) and DALL-E for generating photo-realistic images (Rombach et al., 2022a), WaveNet (Oord et al., 2016) for audio synthesis, and Generative Pre-trained Transformer (GPT; Radford et al.
 2018; 2019, Brown et al.
 2020) for text generation.
-Besides this new wave of generative models, different scientific disciplines have a long history of building data generating models which capture specific processes.
-For example in neuroscience, the occurrence of action potentials is modeled at all different levels of detail (e.g.
-single neuron voltage dynamics (Hodgkin & Huxley, 1952) or at a phenomenological level (Pillow et al., 2008)), whereas in e.g.
-astrophysics there exist various models to simulate galaxy formation (Somerville & Davé, 2015).
-Along generating novel synthetic samples, generative models can be leveraged for specific tasks, such as sample generation conditioned on class labels (e.g., diseased vs.
-healthy brain scans, molecules that can or cannot be synthesized; Urbina et al.
+In addition to the recent surge in generative models, many scientific disciplines have a long history of developing data-generating models that capture specific processes.
+In neuroscience, for example, the occurrence of action potentials is commonly modeled at varying levels of detail (e.g., single neuron voltage dynamics; Hodgkin & Huxley 1952, or at a phenomenological level; Pillow et al.
+2008), whereas in, e.g., astrophysics there exist various models to simulate galaxy formation (Somerville & Davé, 2015).
+Along with generating novel synthetic samples, generative models can be leveraged for specific tasks, such as sample generation conditioned on class labels (e.g., diseased vs.
+healthy brain scans Pombo et al.
+2021, molecules that can or cannot be synthesized; Urbina et al.
 2022, class-conditional image generation; van den Oord et al.
 2016, Dockhorn et al.
-2022, forecasting future states of a dynamical system; Durstewitz et al.
-2023, Jacobs et al.
+2022), forecasting future states of a dynamical system (e.g.,  Jacobs et al.
 2023, Brenner et al.
-2022), data imputation (e.g., Vetter et al.
+2022, Pals et al.
+2024), generating neural population activity conditioned on visual stimuli or behavior (e.g., Molano-Mazon et al.
+2018, Bashiri et al.
+2021, Schulz et al.
+2024, Kapoor et al.
+2024), data imputation (e.g., Vetter et al.
 2023, Lugmayr et al.
 2022), data augmentation for downstream tasks (Rommel et al., 2022), and many more (see also Table S1).
-These powerful capabilities are enabled by the premise that generative models accurately learn a high-dimensional distribution from which we assume our dataset was sampled.
+These powerful capabilities are enabled by the premise that generative models can produce samples from the high-dimensional distribution from which we assume our dataset was sampled.
 The dimensions can correspond to anything from individual pixels or graphs to arbitrary features of physical or abstract objects.
-When aiming to build generative models that better capture the true underlying data distribution, we need to answer a key question: How accurately does our generative model mimic the true data distribution?
+When aiming to build generative models that better capture the true underlying data distribution, we need to answer a key question: How accurately do samples from our generative model mimic those from the true data distribution?
 Manual inspection of generated samples can be a good first check, e.g., in image or audio generation, where we can directly assess the visual likeness or sound quality of the samples (Gerhard et al., 2013, Vallez et al., 2022, Jayasumana et al., 2023).
-In general, however, we would like to have quantitative distances to measure the similarity of distributions, for instance to compare and benchmark different models.
-Many measures have been proposed that provide a quantitative way to assess the similarity of two distributions based on various aspects of their moments, spread, central tendency, and overall probability density (Fig.
+In general, however, we would like to quantitatively compare the similarity of distributions, for instance to benchmark different generative models.
+Many measures have been proposed that assess the similarity of two distributions based on various aspects of their moments or the overall probability density (Fig.
 1).
-Some of these measures require likelihood evaluations, as is possible with generative models such as Gaussian Mixture Models, Normalizing Flows, Variational Autoencoders, autoregressive models or diffusion models (Bishop, 2006, Papamakarios et al., 2021, Box et al., 2015, Kingma & Welling, 2014, Yenduri et al., 2023, Ho et al., 2020, Song et al., 2021).
+Some of these measures require likelihood evaluations, as is possible with generative models such as Gaussian Mixture Models, Normalizing Flows, Variational Autoencoders, autoregressive models or diffusion models (Bishop, 2006, Papamakarios et al., 2021, Box et al., 2015, Kingma & Welling, 2014, Yenduri et al., 2024, Ho et al., 2020, Song et al., 2021).
 However, many contemporary machine learning models (e.g., Generative Adversarial Networks and Energy-Based Models; Goodfellow et al.
@@
 2014, Hinton et al.
-2006) and scientific simulators (e.g., single neuron voltage dynamics; Hodgkin & Huxley 1952) only define the likelihood implicitly, i.e.
-we can not explicitly evaluate their likelihood.
-In this work we focus on distances that can be applied to both classes of models, i.e.
-distances that can be computed only based on generated samples.
+2006) and scientific simulators (e.g., single neuron voltage dynamics; Hodgkin & Huxley 1952) only define the likelihood implicitly, i.e., we can not explicitly evaluate their likelihood.
+Statistical distances that can be computed based on samples only are therefore invaluable for comparing both classes of generative models (and to real data) in scientific contexts.
 Figure 1: The need for statistical distances in scientific generative modeling.
@@
 However, an appropriate sample-based distribution distance $D$ can determine that $p_{2}(x)$ is more similar to $p_{true}(x)$.
-(b) Scientific applications often require evaluating high-dimensional distributions, such as distributions for images or tabular data.
+(b) Scientific applications often require evaluating high-dimensional distributions, such as distributions of images or tabular data.
 In this example, each point represents an X-ray image, where each dimension is one pixel.
-Here, we provide a guide to understanding commonly used sample-based statistical distances.
-Note that with distance, we do not necessarily refer to a distance metric in the mathematical sense (i.e., satisfying symmetry and the triangle inequality) but to a general measure of dissimilarity between two distributions.
-Our goal is not to provide a comprehensive review of statistical distances, as there are already a number of excellent resources for that purpose, especially in specific domains of application (Borji, 2019, Xu et al., 2018, Lopez-Paz & Oquab, 2016, Lueckmann et al., 2021).
-We refer readers to those works for a deeper dive into mathematical properties and empirical comparisons  (Cox et al., 1984, Gibbs & Su, 2002, Basseville, 2013, Cai & Lim, 2022, Muandet et al., 2017, Theis et al., 2016, Betzalel et al., 2022).
-In this guide, we instead focus on four commonly applied sample-based distances in the machine learning literature for evaluating eventually high-dimensional generative models.
-They represent different methodologies for defining statistical distance: Using low-dimensional projections (Sliced-Wasserstein; SW), obtaining a distance using classifiers (Classifier Two-Sample Tests; C2ST), using embeddings through kernels (Maximum Mean Discrepancy; MMD) or neural networks (Fréchet Inception Distance; FID).
-We aim to provide an intuition for how and when to apply these distances, and to build a solid foundation for navigating the extensive literature on statistical distances.
-Towards this goal, we provide for each of the four distances an intuitive and graphical explanation (Section 2).
+Here, we present a guide that aims to serve as an accessible entry point to understanding commonly used sample-based statistical distances.
+Towards this goal, we provide explanations, comparisons, and example applications of four commonly applied distances: Sliced-Wasserstein (SW), Classifier Two-Sample Tests (C2ST), Maximum Mean Discrepancy (MMD), and Fréchet Inception Distance (FID).
+With these resources, we aim to empower researchers to choose, implement, and evaluate the usage and outcomes of statistical distances for generative models in science.
+Note that, with distance, we do not necessarily refer to a distance metric in the mathematical sense (i.e., satisfying symmetry and the triangle inequality) but to a general measure of dissimilarity between two distributions (however, some considered distances are in fact metrics).
+The general and didactic nature of this guide means it can neither be comprehensive nor provide a clear-cut answer as to which distance is ‘best’, as there are numerous choices, and the ‘ideal’ one strongly depends on the domain of application.
+Previous articles have provided useful and extensive reviews for specific use cases.
+For example, Theis et al.
+(2016) discusses commonly used criteria for evaluating generative models of images, Borji (2019), Xu et al.
+(2018), Yang et al.
+(2023) compare metrics specific to evaluating GANs (including, e.g., specialized variants of the FID), Basseville (2013) provides an extensive overview of previous works on divergences, and Gibbs & Su (2002) analyze the theoretical relationships among ‘classic’ distances (including, e.g., Wasserstein and KL-divergence).
+Rather, by going through four different classes of sample-based distances in detail and systematically comparing them on synthetic and real-world applications, we aim to provide a solid foundation for navigating the extensive literature on statistical distances, and to enable readers to reason about other related distances not covered here.
+The outline of this guide is as follows: First, we provide an intuitive and graphical explanation for each of the four distances (Section 2).
 We then perform a systematic evaluation of their robustness as a function of dataset size, data dimensionality, and other factors, such as data multimodality (Section 3).
-Finally, in Section 4, we demonstrate how these distances can be applied to compare generative models in different scientific domains: We evaluate low dimensional models of decision making in behavioral neuroscience and generative models of medical X-ray images.
+Finally, we demonstrate how these distances can be applied to compare generative models in different scientific domains (Section 4): We evaluate low dimensional models of decision making in behavioral neuroscience and generative models of medical X-ray images.
 We show the importance of using multiple complementary distances, as distinct distances can give different results when comparing the same sets of samples.
-By presenting these resources, we aim to empower researchers to choose, implement, and evaluate the use and outcomes of statistical distances for generative models in science.
 2 Sample-based statistical distances
@@
 Computing distance between distributions suffers from the curse of dimensionality, where the computational cost of computing the distance increases very rapidly as the dimensionality of the data increases.
-This problem is especially restricting when the distance is used as part of a loss function in optimization problems, since in this case it needs to be evaluated many times.
+This problem is especially restricting when the distance is used as part of a loss function in optimization problems, since in this case it must be evaluated many times.
 This has prompted the notion of “sliced” distances, which have become increasingly popular in recent years (Kolouri et al., 2019, Nadjahi et al., 2020, Goldfeld & Greenewald, 2021).
-The main idea behind slicing is that for many existing statistical distances we can efficiently evaluate the distance in low-dimensional spaces, especially in one dimension.
+The main idea behind slicing is that for many existing statistical distances we can efficiently evaluate the distance in low-dimensional spaces, especially in a single dimension.
 Therefore, the “slices” are typically one-dimensional lines through the data space (Fig.
@@
 The distance measure of interest between the resulting one-dimensional distributions can then be computed efficiently.
-However, computing the random projection could lead to an unreliable measure of distance: Distinct distributions can produce the same one-dimensional projections.
+However, computing the random projection could lead to an unreliable measure of distance as distinct distributions can produce the same one-dimensional projections.
 Therefore, we repeat the slicing process for many different slices and average the resulting distances.
@@
 2).
-However, we note that slicing has also been done for other distance measures, such as MMD with a specific choice of kernel (Hertrich et al., 2023) and mutual information (Goldfeld & Greenewald, 2021).
-Figure 3: Computing Wasserstein distance.
-Two transport maps mapping the samples from a distribution $p_{1}$ (black) to samples from another distribution $p_{2}$ (blue), shown by arrows.
-The color of the arrow corresponds to the cost (Euclidean distance) between $x_{i}$ and $y_{i}$.
-(a) Randomly chosen transport map.
-(b) The optimal transport map, giving the smallest total cost.
-The total cost for the optimal map in (b) is the Wasserstein distance between these two sets of samples.
-Wasserstein or earth mover distance
+The Wasserstein distance and its sliced variant have several attractive properties: They can be computed differentiably; their computations do not rely heavily on choices of hyperparameters; and the sliced variant is very fast to compute.
+However, we note that slicing has also been done for other distance measures, such as MMD with a specific choice of kernel (Hertrich et al., 2024) and mutual information (Goldfeld & Greenewald, 2021).
+We provide a formal definition of the Wasserstein distance below, and of the Sliced-Wasserstein distance in Appendix A.3.
+Definition of Wasserstein Distance
+We here provide the definition in the common case where probability density functions $p_{1}$ and $p_{2}$ are assumed to exist (formal definition using probability measures in Appendix A.3).
+Let $M\subseteq\mathbb{R}^{d}$, and $||\cdot||_{q}$ be the $q$-norm in $\mathbb{R}^{d}$.
+Then the Wasserstein-$q$ norm can be written as
+$W_{q}(p_{1},p_{2})=\inf_{\gamma\sim\Gamma(p_{1},p_{2})}\left(\mathbb{E}_{x_{1},x_{2}\sim\gamma}\,||x_{1}-x_{2}||_{q}^{q}\right)^{\frac{1}{q}},$ (1)
+where $\Gamma(p_{1},p_{2})$ is the set of all couplings, that is all possible “transportation plans”, between $p_{1}$ and $p_{2}$.
+$\gamma\in\Gamma(p_{1},p_{2})$ is a joint distribution over $(x_{1},x_{2})$ with respective marginals $p_{1}$ and $p_{2}$ over $x_{1}$ and $x_{2}$.
+Sample-Based Wasserstein Distance
+In practice, Eq.
+(1) is analytically solvable for only a few distributions.
+Therefore, Wasserstein distance is typically estimated from finite samples from $p_{1}$ and $p_{2}$.
+However, sample-based estimates of the Wasserstein distance are biased, and the convergence to the true Wasserstein distance is exponentially slower as the dimensionality of the distribution increases (Fournier & Guillin, 2015, Papp & Sherlock, 2022)
 Intuitively, if two given probability distributions are thought of as two piles of dirt, the Wasserstein distance measures the (minimal) cost of “transporting” one pile of dirt to another (Panaretos & Zemel, 2019).
-The formal definition of the Wasserstein distance for continuous distributions, derived from optimal transport, is described in Section A.3.
-Here, we provide a more intuitive definition given a fixed set of samples from two distributions.
+This analogy led to the Wasserstein distance also being known as the Earth Mover’s Distance.
 Suppose we have samples $\{x_{1},\ldots,x_{N}\}\subset\mathbb{R}^{d}$ sampled from a distribution $p_{1}$ and $\{y_{1},...,y_{N}\}\subset\mathbb{R}^{d}$ sampled from another distribution $p_{2}$.
@@
 $C=\begin{bmatrix}D(x_{1},y_{1})&\dots&D(x_{1},y_{N})\\ \vdots&\ddots&\vdots\\ D(x_{N},y_{1})&\dots&D(x_{N},y_{N})\end{bmatrix}$ (2)
-Recalling the earth-mover distance analogy, we want to map each $x_{i}$ to exactly one $y_{j}$, in such a way that the cost of doing so is minimized.
-The minimum transport map is then defining the Wasserstein distance (for the metric $D$) between the two empirical distributions.
-Throughout this work, we use the commonly used Euclidean metric, $L^{2}$, leading to the Wasserstein-2 and Sliced Wasserstein-2 distances.
+Recalling the Earth Mover’s Distance analogy, we want to map each $x_{i}$ to exactly one $y_{j}$, in such a way that the cost of doing so is minimized.
+The minimum transport map then defines the Wasserstein distance (for the metric $D$) between the two empirical distributions.
+Throughout this work, we employ the commonly used Euclidean metric, $L^{2}$, leading to the Wasserstein-2 and Sliced Wasserstein-2 distances.
 More precisely, we define a “transport map” to be a permutation matrix, $\pi\in\{0,1\}^{N\times N}$, which is a matrix with exactly one nonzero entry in each row.
 The entry $\pi_{ij}=1$ means that we transport the point $x_{i}$ to the point $y_{j}$.
-Then finding the transport map that minimizes the overall cost can be stated as
+Then, finding the transport map that minimizes the overall cost can be stated as
 $\pi^{*}=\min_{\pi}\sum_{ij}\pi_{ij}C_{ij}.$ (3)
@@
 3b.
+Figure 3: Computing Wasserstein distance.
+Two transport maps mapping the samples from a two-dimensional distribution $p_{1}$ (black) to samples from another distribution $p_{2}$ (blue), shown by arrows.
+The color of the arrow corresponds to the cost (Euclidean distance) between $x_{i}$ and $y_{i}$.
+(a) Randomly chosen transport map.
+(b) The optimal transport map, giving the smallest total cost.
+The total cost for the optimal map in (b) is the Wasserstein distance between these two sets of samples.
+Note that this schematic demonstrates transport maps for the non-sliced Wasserstein distance in two dimensions.
 Slicing Wasserstein brings efficiency
 Solving the optimal transport problem (Eq.
-(2)) with the Hungarian method has a time complexity of $O(N^{3})$ in the number of samples $N$ (although faster approximations exists, see Peyré et al.
+(3)) with the Hungarian method has a time complexity of $O(N^{3})$ in the number of samples $N$ (although faster $O(N^{2}\log N$) approximations exist, see Peyré et al.
 2017).
 However, in the special case where the data is one-dimensional, the Wasserstein distance can be calculated by sorting the two datasets, obtaining the order statistics $\{x_{(1)},..,x_{(N)}\}$ and $\{y_{(1)},...,y_{(N)}\}$ and computing the sum of the distances $\sum_{i}D(x_{(i)},y_{(i)})$.
-This has a time complexity of $O(N\log(N))$.
+This has a time complexity of $O(N\log N)$.
 Thus, slicing the Wasserstein distance with one-dimensional projections becomes very efficient.
-While the value of the SW distance does not converge to the Wasserstein distance, even in the case of an infinite number of data samples and slices, the SW distance is a metric (in the mathematical sense) as long as $D$ is a metric on $\mathbb{R}^{d}$ and it acts as a lower bound to the Wasserstein distance (Nadjahi, 2021).
-The Wasserstein distance and its sliced variant have several attractive properties: they can be computed differentiably; their computations do not rely heavily on choices of hyperparameters; and the sliced variant is very fast to compute.
-However, a disadvantage of the Wasserstein distance is its transparency: The numerical value of the Wasserstein distance has no intuitive interpretation due to its definition in terms of optimal transport maps.
-Additionally, per definition, the sliced variant is insensitive to differences within orthogonal subspaces of the slices.
-To still capture differences in all dimensions, naively, one would have to increase the number of slices (in the worst case) exponentially with the dimension, diminishing computational efficiency.
-However, other approaches exist to reduce this problem for nonlinear (Kolouri et al., 2019) or other specific (Deshpande et al., 2019; 2018) slices.
-Furthermore, slicing may also be relaxed to other kinds of data specific projections, such as Fourier features for stationary time series or locally connected projections for images (Du et al., 2023, Cazelles et al., 2020).
+While the value of the SW distance does not converge to the Wasserstein distance, even in the case of an infinite number of data samples and slices, the SW distance is a metric (in the mathematical sense) as long as $D$ is a metric on $\mathbb{R}^{d}$, and it acts as a lower bound to the Wasserstein distance (Nadjahi, 2021).
+Limitations
+When two high-dimensional distributions differ only along a small subset of directions, relying on one dimensional projections can become a significant limitation.
+This is because in high-dimensional spaces, most slices will be (almost) orthogonal to the few directions which distinguish the two distributions.
+In such a situation, the gain in efficiency from slicing can be impeded by the large number of slices we would need to reliably tell the two distributions apart.
+This limitation can be addressed by considering nonlinear (Kolouri et al., 2019) or other specific (Deshpande et al., 2019; 2018) slices.
+Furthermore, slicing may also be relaxed to other kinds of data-specific projections, such as Fourier features for stationary time series or locally-connected projections for images (Du et al., 2023, Cazelles et al., 2021).
+Finally, while sliced distances can overcome computational challenges of computing distances in higher dimensions, they typically inherit the limitations of the distance being sliced.
 2.2 Classifier-based: Classifier Two-Sample Test (C2ST)
@@
 The Classifier Two-Sample Test (C2ST) uses a classifier that discriminates between samples from two distributions (Fig.
-4a) (Lopez-Paz & Oquab, 2016, Friedman, 2003).
-The distance between the distributions can then be quantified with various measures of classifier performance.
+4a;  Lopez-Paz & Oquab 2017, Friedman 2003).
+The distance between the distributions can then be quantified by various measures of classifier performance.
+For example, one would train a classifier $c(x)$ to distinguish samples from the generative model and the data, and then evaluate the C2ST as $\frac{1}{2}[\mathbb{E}_{p(x)}[\mathds{1}(c(x)=0)]+\mathbb{E}_{q(x)}[\mathds{1}(c(x)=1)]]$.
 The classification accuracy provides a particularly intuitive and interpretable measure of the similarity of the distributions.
@@
 C2ST can also be used to test the statistical significance of the difference between two sets of samples.
-Unlike other measures, however, C2ST can be expensive to compute because it requires training a classifier and using the classification accuracy as a differentiable training objective is not straightforward (see Section A.5.1).
+However, calculating C2ST can be resource-intensive because it involves training a classifier.
+Additionally, using classifiers in a differentiable training objective can be challenging (but is sometimes possible, as in, e.g., Generative Adversarial Networks;  Goodfellow et al.
+2014).
 Furthermore, the value is dependent on the capacity of the classifier, and hence on many hyperparameters such as classifier architecture or training procedure.
-In our experiments, we used a scikit-learn Multi-Layer-Perceptron classifier, combined with a five-fold cross-validation routine to estimate the accuracy returned (Pedregosa et al., 2011).
+This dependence on a trained classifier can result in C2ST estimates that are biased, and the variety of possible classifier architectures means theoretical guarantees such as sample complexity are difficult to determine.
+In our experiments, we used a scikit-learn Multi-Layer Perceptron classifier, combined with a five-fold cross-validation routine to estimate the accuracy returned (Pedregosa et al., 2011).
 Common failure modes
-As mentioned above, for any realistic scenario, the C2ST is computed by training a classifier.
+As mentioned above, for any realistic scenario the C2ST is computed by training a classifier.
 The resulting C2ST will only be a good measure of distance between real and generated data if the classifier is close or equal to the optimal classifier.
@@
 If the C2ST is estimated with a neural network, however, we observe that this C2ST can be systematically underestimated: for example, when only few samples from the data and generative model are available, the neural network predicts a C2ST of 0.5 (Fig.
-5a, top right) – in other words, it predicts that the generative model and the data follow the same distribution.
-Similarly, if the neural network is not expressive enough, e.g.
-with too few hidden units, the classifier will return a low C2ST, around 0.5 (Fig.
+5a, top right)—in other words, it predicts that the generative model and the data follow the same distribution.
+Similarly, if the neural network is not expressive enough, e.g., with too few hidden units, the classifier will also return a low C2ST, around 0.5 (Fig.
 5a, bottom right).
-These issues can make the C2ST easy to misuse: In many cases, reporting a low C2ST is desirable for generative models since it indicates that the model perfectly matches data, but one can achieve a low C2ST simply by not investing sufficient time into obtaining a strong classifier.
+These issues can make the C2ST easy to misuse.
+In many cases, reporting a low C2ST is desirable for generative models since it indicates that the model perfectly matches data, but one can achieve a low C2ST simply by not investing sufficient time into obtaining a strong classifier.
 C2ST can remain very high even for seemingly good generative models
-We previously argued that the C2ST is an interpretable measure – while this is generally true, the C2ST can sometimes be surprisingly high even if the generative model seems well aligned with the data.
+We previously argued that the C2ST is an interpretable measure—while this is generally true, the C2ST can sometimes be surprisingly high even if the generative model seems well aligned with the data.
 For example, when the generative model aligns very well with the data for every marginal, the C2ST can still be high if the data is high-dimensional (Fig.
@@
 Other C2ST variants
-While we focus on a standard C2ST definition by using classification accuracy as the C2ST distance (Lopez-Paz & Oquab, 2016), any other performance metric for binary classification could be used (Raschka, 2014).
+While we focus on a standard C2ST definition by using classification accuracy as the C2ST distance (Lopez-Paz & Oquab, 2017), any other performance measure for binary classification could be used (Raschka, 2014).
 Kim et al.
-(2019) even argue that classic accuracy is sub-optimal due to the “binarization” of the class probabilities and proceeds to instead use the mean squared error between the predicted and ‘target’ value of 0.5.
-Other approaches instead construct a likelihood ratio statistic (Pandeva et al., 2022).
+(2019) argues that classic accuracy is sub-optimal due to the “binarization” of the class probabilities and proceeds to instead use the mean squared error between the predicted and “target” value of 0.5.
+Other approaches instead construct a likelihood ratio statistic (Pandeva et al., 2024).
 Additionally, instead of using the estimated class probabilities, Cheng & Cloninger (2022) consider using the average difference in logits (i.e., activations in the last hidden layer).
-However, classification accuracy is still the most commonly used variant of C2ST.
+We note that the learned classifier in C2ST can be applied to estimations of a density ratio $\frac{p(x)}{q(x)}$, also known as the likelihood ratio trick (Hastie et al., 2001, Sugiyama et al., 2012).
+Density ratio estimation has attracted a great deal of attention in the statistics and machine learning communities since it can be employed for estimating divergences between two distributions, such as the Kullback–Leibler divergence (Titsias & Ruiz, 2019, Huszár, 2017) and Pearson divergence (Srivastava et al., 2020).
 2.3 Kernel-based: maximum mean discrepancy (MMD)
-Figure 6: Maximum mean discrepancy (MMD).
-(a) Two example distributions $p_{1}(x)$, $p_{2}(x)$ and observed data $p_{true}(x)$ that we want to compare.
-(b) MMD can be defined as the difference between the expectations of some embedding function $\phi(x)$.
-If we take the identity as embedding ($\phi^{(1)}(x)=x$; left), we end up computing the differences between the means of the distributions, which are all equal for the three distributions.
-If we add a quadratic feature ($\phi(x)^{(2)}=[x,x^{2}]^{\mathsf{T}}$; right), we can distinguish distributions with different variances.
-Note that we still have $\mathsf{MMD}^{2}[\phi^{(2)},p_{2},p_{true}]=0$, despite $p_{2}$ being different from $p_{true}$ (c) Using the kernel trick we can avoid computing the embeddings all together but use implicit embeddings that capture all relevant features of the distributions.
-MMD is a popular distance metric that is applicable to a variety of data domains, including high-dimensional continuous data spaces, strings of text as well as graphs (Borgwardt et al., 2006, Gretton et al., 2012a, Muandet et al., 2017).
+MMD is a popular distance metric that is applicable to a variety of data domains, including high-dimensional continuous data spaces, strings of text, and graphs (Borgwardt et al., 2006, Gretton et al., 2012a, Muandet et al., 2017).
 It has been used to evaluate generative models (Sutherland et al., 2021, Borji, 2019, Lueckmann et al., 2021) and also has the ability to indicate where the model and the true distribution differ (Lloyd & Ghahramani, 2015).
 The distance provided by MMD can straightforwardly be used to test whether the difference between two sets of high-dimensional samples is statistically significant (Gretton et al., 2012a).
-To assess whether two set of samples are drawn from the same distribution, MMD makes use of a kernel function to (implicitly) embed the samples via an embedding function $\phi$, also called feature map.
-If we choose the right kernel, we can end up embedding our samples in a space where properties of the underlying distributions are easily compared.
+To assess whether two sets of samples are drawn from the same distribution, MMD makes use of a kernel function to (implicitly) embed the samples via an embedding function $\phi$, also called a feature map.
+If we choose the right kernel, we can end up embedding our samples in a space where the properties of the underlying distributions are easily compared.
 We will motivate the use of the kernel in MMD by illustrating different explicit embeddings before introducing the implicit embedding via a kernel $k$.
@@
 for any embedding function $\phi$.
-If we want to compare samples of real numbers from two distributions $p_{1}$ and $p_{2}$ (Fig.
-6a), we can think about different embedding functions $\phi$ to compare these.
+Figure 6: Maximum mean discrepancy (MMD).
+(a) Two example distributions $p_{1}(x)$, $p_{2}(x)$ and observed data $p_{true}(x)$ that we want to compare.
+(b) MMD can be defined as the difference between the expectations of some embedding function $\phi(x)$.
+If we take the identity as embedding ($\phi^{(1)}(x)=x$; left), we end up computing the differences between the means of the distributions, which are all equal for the three distributions.
+If we add a quadratic feature ($\phi^{(2)}(x)=[x,x^{2}]^{\mathsf{T}}$; right), we can distinguish distributions with different variances.
+Note that we still have $\mathsf{MMD}^{2}[\phi^{(2)},p_{2},p_{true}]=0$, despite $p_{2}$ being different from $p_{true}$ (c) Using the kernel trick we can avoid computing the embeddings all together but use implicit embeddings that capture all relevant features of the distributions.
+If we have samples of real numbers from two distributions $p_{1}$ and $p_{2}$ (Fig.
+6a), there are various embedding functions $\phi$ we could use to compare these.
 The simplest possible function $\phi^{(1)}:\mathbb{R}\rightarrow\mathbb{R}$ is the identity mapping $\phi^{(1)}(x)=x$ (Fig 6b, left).
-However, in this case the MMD will simply be the absolute difference between the means (first moments) of the distributions (for details, see Section A.2):
+However, in this case, the MMD will simply be the absolute difference between the means (first moments) of the distributions (for details, see Section A.2):
 $\mathsf{MMD}[\phi^{(1)},p_{1},p_{2}]$ $=\lvert\mu_{p_{1}}-\mu_{p_{2}}\rvert.$
-As both models and the true distribution in Fig.
-6 have the same mean, this does not yet let us discriminate between them.
+This does not yet allow us to discriminate different distributions with equal means (all distributions in Fig.
+6).
 If we now expand our embedding with a quadratic term, $\phi^{(2)}:\mathbb{R}\rightarrow\mathbb{R}^{2}$ as $\phi^{(2)}(x)=\begin{bmatrix}x\\ x^{2}\end{bmatrix}$ (Fig 6b, right), the MMD yields (for details, see Section A.2)
@@
 We can now rewrite the inner product $\langle\phi(x),\phi(x^{\prime})\rangle$ in terms of a kernel function $k$: $\langle\phi(x),\phi(x^{\prime})\rangle=k(x,x^{\prime})$.
-Thus, if we can find a kernel for our feature map, we can avoid explicitly computing the features altogether but instead, we directly compute
+Thus, if we can find a kernel for our feature map, we can avoid explicitly computing the features altogether and instead directly compute
 $\mathsf{MMD}^{2}[k,p_{1},p_{2}]=\mathbb{E}_{p_{1}(x),p_{1}^{\prime}(x^{\prime})}[k(x,x^{\prime})]+\mathbb{E}_{p_{2}(y),p_{2}^{\prime}(y^{\prime})}[k(y,y^{\prime})]-2\mathbb{E}_{p_{1}(x),p_{2}(y)}[k(x,y)].$
@@
 6c).
-If we can define a kernel whose corresponding embedding captures all, potentially infinitely many moments, we would have an MMD that is zero only if two distributions are exactly equal.
-These kernels are called characteristic (Section A.2,  Gretton et al.
-2012a), and include the commonly used Gaussian kernel: $k_{G}(x,x^{\prime})=\exp(-\frac{\lVert x-x^{\prime}\rVert^{2}}{2\sigma^{2}})$ (see, e.g., Sriperumbudur et al.
-(2009) for other characteristic kernel choices).
+If we can define a kernel whose corresponding embedding captures all, potentially infinitely many moments, we would have an MMD that is zero only if two distributions are exactly equal and the MMD becomes a metric.
+These kernels are called characteristic (Section A.2Gretton et al.
+2012a), and include the commonly used Gaussian kernel: $k_{G}(x,x^{\prime})=\exp(-\frac{\lVert x-x^{\prime}\rVert^{2}}{2\sigma^{2}})$.
+A number of other kernels can also be used (see, e.g., Sriperumbudur et al.
+2009).
+For instance, using a kernel induced by the Euclidean distance, MMD can be shown to be equivalent to the standard energy distance (Székely & Rizzo, 2013).
+In fact, a wider equivalence between MMD and the generalized energy distance has been established, using distance-induced kernels (Sejdinovic et al., 2013).
 MMD in practice
-Typically the kernel version of MMD is used, which is straightforwardly estimated with its empirical (unbiased) estimate:
+Typically, the kernel version of MMD is used, which is straightforwardly estimated with its empirical, unbiased, estimate:
 $\mathsf{MMD}^{2}=\frac{1}{m(m-1)}\sum_{i}^{m}\sum_{j\neq i}k(x_{i},x_{j})+\frac{1}{n(n-1)}\sum_{i}^{n}\sum_{j\neq i}k(y_{i},y_{j})-\frac{2}{mn}\sum_{i,j}k(x_{i},y_{j}).$
-MMD in this form can be applied to many forms of data, as long as we can define a kernel, which can include graphs (Vishwanathan et al., 2010, Gärtner, 2003) or strings of text (Lodhi et al., 2002), in addition to vectors and matrices.
+MMD in this form can be applied to many forms of data as long as we can define a kernel, which can include graphs (Vishwanathan et al., 2010, Gärtner, 2003) or strings of text (Lodhi et al., 2002), in addition to vectors and matrices.
 When we estimate the MMD with a finite number of samples, the selection of the right kernel and its parameters becomes crucial.
 For example, when using a Gaussian kernel, one has to choose the bandwidth $\sigma$.
-The MMD approaches zero if we take $\sigma$ to be close to zero (then $k_{G}(x,x^{\prime})=1\;\mathsf{if}\;x=x^{\prime}\;\mathsf{else}\;k_{G}(x,x^{\prime})\rightarrow 0$) or if $\sigma$ is large (then $k_{G}(x,x^{\prime})\rightarrow 1\;\forall\;x,x^{\prime}$(Gretton et al., 2012a).
+The MMD approaches zero if we take $\sigma$ to be close to zero (then $k_{G}(x,x^{\prime})=1\;\mathsf{if}\;x=x^{\prime}\;\mathsf{else}\;k_{G}(x,x^{\prime})\rightarrow 0$) or if $\sigma$ is large (then $k_{G}(x,x^{\prime})\rightarrow 1\;\forall\;x,x^{\prime}$; Gretton et al.
+2012a).
 A common heuristic to remedy this parameter choice is picking the bandwidth based on the scale of the data.
-The median heuristic set the bandwidth to the median distance between points in the aggregate sample (Gretton et al., 2012a).
+The median heuristic sets the bandwidth to the median distance between points in the aggregate sample (Gretton et al., 2012a).
 Another common approach is based on cross-validation, or data splitting (Gretton et al., 2012a; b, Jitkrittum et al., 2016, Sutherland et al., 2021): The dataset is divided, with a hold-out set used for kernel selection, and the other part used for evaluating MMD.
 While the data splitting method does not involve any heuristic, it can lead to errors in MMD since it reduces the number of data points available for estimating the MMD.
-Recent work attempts to choose hyperparameters without employing data splitting or any heuristic (Biggs et al., 2023, Schrab et al., 2023, Kübler et al., 2022b; a).
-While we aim in general for a kernel that captures the (dis)similarity between the data points well, such a kernel can be domain specific or specifically designed for downstream analysis tasks.
-The similarity between two strings (e.g., DNA sequences, text) can for instance be estimated by looking at the frequency of small subsequences (Leslie et al., 2001, Lodhi et al., 2002).
-It is furthermore possible to aggregate simpler kernels into a more expressive one (Gretton et al., 2012b), or to use a deep kernel (i.e., based on neural networks) that can exploit features of particular data modality such as images (Liu et al., 2020, Gao et al., 2021).
+Recent work attempts to choose hyperparameters without employing data splitting (Biggs et al., 2023, Schrab et al., 2023, Kübler et al., 2022b; a).
+We often aim for a kernel that captures the (dis)similarity between the data points well, such a kernel can be domain-specific or specifically designed for downstream analysis tasks.
+The similarity between two strings (e.g., DNA sequences or text) can, for instance, be estimated by looking at the frequency of small subsequences (Leslie et al., 2001, Lodhi et al., 2002).
+Furthermore, it is possible to aggregate simpler kernels into a more expressive one (Gretton et al., 2012b), or to use a deep kernel (i.e., based on neural networks) that can exploit features of particular data modalities such as images (Liu et al., 2020, Gao et al., 2021).
 2.4 Network-based: Embedding-space measures
-Distribution comparisons on structured data spaces, such as the set of natural images, present unique challenges.
+Distribution comparisons on structured data spaces, such as a set of natural images, present unique challenges.
 Such data is usually high-dimensional (high-resolution images) and contains localized correlations.
 Furthermore, images of different object classes (such as airplanes and dogs) share low-level features in the form of edges and textural details but differ in semantic meaning.
-Similar challenges occur for time-series data, natural language text, and other complex data type (Smith & Smith, 2020, Jeha et al., 2021).
+Similar challenges occur for time-series data, natural language text, and other complex data types (Smith & Smith, 2020, Jeha et al., 2021).
 Figure 7: Network-based metrics.
-Instead of directly computing distances in data space, complex data e.g., natural images of dogs sampled from $p_{1}(x)$ and aircraft sampled from $p_{2}(y)$, are jointly embedded into a vector space.
+Instead of directly computing distances in data space, complex data, e.g., natural images of dogs sampled from $p_{1}(x)$ and aircraft sampled from $p_{2}(y)$, are jointly embedded into a vector space.
 The embedding function can, for example, be a deep neural network.
 The resulting distributions in feature space are then compared by a classical measure of choice $D$.
-In this section, we take the example of natural images, but the presented framework generalizes to other data types.
+In this section, we rely on the example of natural images, but the presented framework generalizes to other data types.
 Naive distances would operate on a per-pixel basis, leading to scenarios where, for example, white dogs and black dogs are considered vastly different despite both being categorized as dogs.
-As we would like to have a distance measure that operates based on details relevant to the comparison we can leverage neural networks trained on a large image dataset that captures features ranging from low-level to high-level semantic details: While earlier layers in a convolutional neural network focus on edge detection, color comparison, and texture detection, later layers learn to detect high-level features, such as a dog’s nose or the wing of an airplane, which thought to be relevant for a meaningful comparison.
+As we would like to have a distance measure that operates based on details relevant to the comparison, we can leverage neural networks trained on a large image dataset that captures features ranging from low-level to high-level semantic details: While earlier layers in a convolutional neural network focus on edge detection, color comparison, and texture detection, later layers learn to detect high-level features, such as a dog’s nose or the wing of an airplane, which thought to be relevant for a meaningful comparison.
 Embedding-based distances use these activations of neural network layers as an embedding to compare the image distributions.
-The most popular distance in this class is the Fréchet Inception Distance (FID) (Heusel et al., 2017), used to evaluate generative models for images.
-The FID uses a convolutional neural net’s embeddings (specifically InceptionV3 (Szegedy et al., 2015a)) to extract the relevant features, applies a Gaussian approximation in the embedding space, and computes the Wasserstein distance on this approximation.
-An FID-like measure, in essence, requires a suitable embedding network $f:\mathcal{X}\rightarrow\mathbb{R}^{d}$, where $f$ transforms the data from the original high-dimensional space $\mathcal{X}$ into a lower-dimensional, feature-rich representation in $\mathbb{R}^{d}$ (Fig.
+The most popular distance in this class is the Fréchet Inception Distance (FID; Heusel et al.
+2017), used to evaluate generative models for images.
+The FID uses a convolutional neural network’s embeddings (specifically InceptionV3 (Szegedy et al., 2015)) to extract relevant features, applies a Gaussian approximation in the embedding space, and computes the Wasserstein distance on this approximation.
+A FID-like measure, in essence, requires a suitable embedding network $f:\mathcal{X}\rightarrow\mathbb{R}^{d}$, where $f$ transforms the data from the original high-dimensional space $\mathcal{X}$ into a lower-dimensional, feature-rich representation in $\mathbb{R}^{d}$ (Fig.
 7).
@@
 Under this Gaussian approximation, the squared Wasserstein distance (also known as the Fréchet distance) can be analytically computed as
-$W^{2}((\mu_{1},\Sigma_{1}),(\mu_{2},\Sigma_{2}))=\|\mu_{1}-\mu_{2}\|^{2}+\text{Tr}\left(\Sigma_{1}+\Sigma_{2}-2\left(\Sigma_{1}\Sigma_{2}\right)^{\frac{1}{2}}\right).$
+$W^{2}((\mu_{1},\Sigma_{1}),(\mu_{2},\Sigma_{2}))=\|\mu_{1}-\mu_{2}\|^{2}+\text{Tr}\left(\Sigma_{1}+\Sigma_{2}-2\left(\Sigma_{1}\Sigma_{2}\right)^{\frac{1}{2}}\right).$ (4)
 In principle, any appropriate metric can be used in place of the Fréchet distance.
-Jayasumana et al.
-(2023) show that the MMD (Section 2.3) can be better suited as a metric in the embedding space, as it perceptually matches human judgement on assessing image quality and coverage in generative models.
+For instance, Jayasumana et al.
+(2023), Bińkowski et al.
+(2018), Xu et al.
+(2018) use MMD as a metric in the embedding space, and the MMD-based Inception Distance is often referred to as Kernel Inception Distance (KID).
+KID is known to have some advantages over FID: unlike FID, KID has a simple, unbiased estimator with better sample efficiency (Jayasumana et al., 2023), and does not assume any parametric forms for the distributions.
+Since KID involves MMD, we must carefully select the proper kernel and its hyperparameter when applying it.
 A related and commonly used quality measure for images is the Inception Score (Salimans et al., 2016).
 In contrast to the FID, this measure uses the average InceptionV3 predicted class probabilities and compares them with the true marginal class distribution.
-Note that while both this score and the FID can agree with traditional distances (e.g.
-certain divergences), they might evaluate models differently (Betzalel et al., 2022); see Barratt & Sharma (2018) for further limitations of the Inception Score.
+Note that while both this score and the FID can agree with traditional distances (e.g., certain divergences), they might evaluate models differently (Betzalel et al., 2022); see Barratt & Sharma (2018) for further limitations of the Inception Score.
 Limitations
-One of the biggest limitations is the requirement of a suitable embedding net.
-Newer and more robust networks, such as the image network of the CLIP (Radford et al., 2021) vision-language model, provide better and semantically more consistent embeddings (Betzalel et al., 2022) than the InceptionV3 network.
+One of the biggest limitations is the requirement of a suitable embedding network.
+Newer and more robust networks, such as the image network of the CLIP (Radford et al., 2021) vision-language model, provide better and more semantically consistent embeddings (Betzalel et al., 2022, Jayasumana et al., 2023) than the InceptionV3 network.
 However, as the embedding network is generally non-injective, identical distributions in the embedding space may not necessarily translate to identical distributions in the original space.
@@
 2022).
-3 Comparison and scalability
-When evaluating (or training) generative models, it is important to understand that different statistical distances penalize different features of the generated samples.
-They might for instance weigh differently how important it is to have large sample variability versus how well the modes of the true distribution are captured (as can be illustrated by optimizing a mis-specified model using the different distances, see Fig.
-S1 and  Theis et al.
+3 Comparing distances: sample size, dimensionality, and more
+In this section, we empirically compare these distances along several dimensions, with additional results presented in the Appendices.
+When evaluating (or training) generative models, it is important to understand that different statistical distances pay attention to different features of the generated samples.
+A key aspect to consider is that differences between distances can be especially pronounced in applications where we only have a limited amount of data points, e.g., identifying rare cell types (Marouf et al., 2020), or where we have very high-dimensional data, e.g., neural population recordings in neuroscience (Stringer et al., 2019).
+In such cases, one needs to ensure that the distance measures can reliably distinguish different distributions for the given sample set size while remaining computationally tractable.
+We therefore perform a number of empirical studies here: First, we investigate SW, C2ST and MMD when it comes to distinguishing data sets given varying numbers of samples and data with varying dimensionality (Section 3.1).
+Second, we investigate the scaling properties of FID on the ImageNet dataset (Section 3.2), as well as considering the other distances when computed on the InceptionV3 image embeddings.
+We also summarize the theoretical sample and computational complexities of the four discussed distances in Section A.6.
+Furthermore, different distances might weigh differently the importance of sample diversity, or mode coverage, versus how well the modes of the true distribution are captured.
+If models only capture one mode, or a subset of the modes, i.e., exhibit mode collapse, they might have learned to generate realistic but unvaried samples.
+We illustrated the trade-off between mode collapse and mode coverage by optimizing mis-specified (Fig.
+S2) and well-specified (Fig.
+S3) models using the different distances (see also Section A.7 and Theis et al.
 2016).
-Thus, using different statistical distances to evaluate the quality of generative models can lead to different conclusions.
 Figure 8: Scalability of different statistical distances with sample size and dimensionality.
-(a,b) Comparison of sample sets with varying sample size (between 50 and 4k samples per set) of a ‘true’ distribution either with a second dataset of the same distribution or with a sample set from an approximated/shifted distribution.
+(a,b) Comparison of sample sets with varying sample size (between 50 and 4k samples per set) of a ‘true’ distribution, either with a second set of samples of the same distribution or with a sample set from a model consisting of an approximated/shifted distribution.
 We show the mean and standard deviation over five runs of randomly sampled data.
@@
 (a) Distances for the 2d-MoG example shown in Fig.
-1 compared to samples from a unimodal Gaussian approximation with the same mean and covariance.
-(b) Distances for a ten-dimensional standard normal distribution, for which the first dimension is shifted by one for the shifted dataset.
+1 compared to samples from a uni-modal Gaussian approximation with the same mean and covariance.
+(b) Distances for a ten-dimensional standard normal distribution, with a model whose first dimension is shifted by one.
 (c) Distances based on 10k samples from a standard normal distribution with varying dimensions (between 5 and 1000).
-As in (b), the first dimension is shifted by one for the ‘shifted’ dataset.
+As in (b), the first dimension of the model is shifted by one.
 We show the mean and standard deviation over five runs of randomly sampled data.
 One MMD bandwidth was selected for all n-dimensional datasets.
-These differences can be even more pronounced in applications where we only have a limited amount of data points, e.g., identifying rare cell types (Marouf et al., 2020), or where we have very high-dimensional data, e.g., neural population recordings in neuroscience (Stringer et al., 2019).
-In such cases, one needs to ensure that the distance measures can reliably distinguish different distributions for the given sample set size while remaining computationally tractable.
-In the following sections, we investigate the sensitivity of the three presented distances which do not rely on embeddings, when it comes to distinguishing data sets with varying numbers of samples (Section 3.1) and varying numbers of dimensions (Section 3.2).
-As the absolute values of the distance measures are often hard to interpret and different measures are on different scales, we examined the relative distances by comparing two or more models to the true data.
-We, therefore, applied the distances to compare samples from a ‘true’ distribution against itself (intra-dataset) and against samples from another distribution that is either an approximation or a shifted version of the true distribution (inter-dataset).
-In our experiments, we focused on inter- and intra-dataset comparisons for the following three datasets: First, we compared the two-dimensional Mixture of Gaussians (“2d-MoG”) dataset introduced in Fig.
-1 with samples from a unimodal Gaussian approximation with the same mean and covariance as the true data samples.
-Second, we compared samples from a ten-dimensional standard normal distribution with a shifted normal distribution for which the first dimension is shifted by one (“10-dim Gaussian”).
-And last, we compared a standard normal distribution with varying dimensionality to shifted distributions for which we respectively shifted the first dimension for the inter-dataset comparison (“n-dim Gaussian”).
+Note that the absolute values of the distance measures can be hard to interpret and different measures lie on different scales.
+For this reason, we compare the distance between a set of sample from the model and a set of samples from the true distribution to the distance between two sets of samples from the true distribution.
+3.1 Varying number of samples and dimensionality
+Here, we assessed SW, C2ST and MMD using three datasets: First, we compared the two-dimensional Mixture of Gaussians (“2d-MoG”) dataset introduced in Fig.
+1 with samples from a model consisting of a unimodal Gaussian approximation with the same mean and covariance as the true data.
+Second, we compared samples from a ten-dimensional standard normal distribution with a model that is also a normal distribution that matches the true distribution in all but the first dimension—where the mean is shifted by one (“10-dim Gaussian”).
+Last, we repeated the second experiment, now taking a normal distribution with varying dimensionality as true dataset, and a model who again matches in all dimensions, except for the shifted first dimension (“n-dim Gaussian”).
 We used default parameters for the SW and C2ST measures while adjusting the bandwidth parameter for the MMD measure for each of the three comparisons.
-Finally, as FID and other network-based distances do require an embedding network, we investigate the scaling properties specific to FID on the ImageNet dataset (in Section 3.3).
-3.1 Varying number of samples
+Sample size
 We explored the robustness of the distances to low sample sizes on the 2d-MoG and the 10-dim Gaussian dataset.
-We found that for the 2d-MoG dataset, all measures failed to reflect the dissimilarity of the distributions at the lowest sample size of 50 samples (Fig.
+We found that for the 2d-MoG dataset, computed values quickly stabilized for all measures by a sample size of 1000 and yielded the expected results of indicating low differences between two sets of samples from the true distribution (true), and a higher difference between samples from the true and model distributions (approx.).
+All measures failed to reflect the dissimilarity of the distributions at the lowest sample size of 50 samples (Fig.
 8a).
-However, C2ST’s behavior differs from MMD and SW, with C2ST indicating that the distributions are similar (C2ST $\approx$ 0.5) for both intra- (true) and inter-dataset (approx.) comparisons while the other two distances indicate they are different (distance $\neq 0$).
-The malfunction of C2ST can be harder to detect in such cases, compared to the one of MMD and SW.
-While the latter is easily identified by the incorrect intra-dataset results, the malfunction of C2ST is hard to detect for unknown distributions.
-For all measures, computed values quickly stabilized by a sample size of 1000 and yielded the expected results of low intra-dataset differences and high inter-dataset differences.
-For the 10-dim Gaussian dataset, we observed that all distances can identify samples from the same distribution as more similar than samples from different distributions (Fig.
-8b); while all three distances struggle with low sample sizes (see also Supp.
+Here C2ST’s behavior differs from MMD and SW, with C2ST indicating that the distributions are similar (C2ST $\approx$ 0.5) while the other two distances indicate they are different (distance $\neq 0$).
+For the 10-dim Gaussian dataset, we observed that all three distances can robustly identify samples from the same distribution as more similar than samples from different distributions with no measure being clearly superior to the others (Fig.
+8b).
+For the previous 2d-MoG experiment, more samples were required to clearly detect the difference between the two distributions (Fig.
+8a) as compared to the 10-dim Gaussian (Fig.
+8b).
+Intuitively, the more pronounced the differences in the distributions we compare, the fewer samples we need to detect these differences (see additional experiments in Supp.
 Fig.
-S3, given enough samples, they all become robust, with no measure being clearly superior to the others.
-For the 2d-MoG experiment, more samples are required to clearly detect the difference between the two distributions (Fig.
-8a) as compared to the case where the mean in one dimension is shifted (Fig.
-8b).
-Intuitively, the larger the differences in the distributions we want to compare, the fewer samples we need to detect these differences (see additional experiments Supp.
-Fig.
-S3).
-3.2 Varying dimensionality
-We further tested how the distances scale with the data dimension using $n$-dimensional standard normal distributions.
-In the first experiment, the shifted distribution differed only in the first dimension, which was mean-shifted by one.
-As the dimensionality increases, all distances reliably indicate no difference in intra-dataset comparisons , but C2ST is the only measure that consistently identifies the inter-dataset difference (Fig.
+S4).
+Dimensionality
+We further tested how the distances scale with the data dimension using the n-dim Gaussian dataset.
+As the dimensionality increases, all distances indicate no difference between two sets of samples from the true distribution, except C2ST which is the only measure that consistently identifies samples from the true and model as distinct (Fig.
 8c).
-Note that this dataset is different from Fig.
-5b, where mean shifts were applied to every dimension.
 When we changed the structure of the data distribution (e.g., by changing the mean of all dimensions or their variances, see Supp.
 Fig.
-S4), we observed a similar picture with some particularities: While the SW distance with a fixed number of slices has difficulties if the disparity between the distribution is only in one dimension, its performance drastically improves for differences in all dimensions, which is expected from the random projections SW is performing.
+S5), we observed a similar picture with some particularities: While the SW distance with a fixed number of slices has difficulties if the disparity between the distribution is only in one dimension, its performance drastically improves for differences in all dimensions, which is expected from the random projections SW is performing.
 C2ST seems to robustly detect differences even in high dimensions in these modified datasets, though previous experiments showed that this measure can be oversensitive to small changes in high dimensions (Fig.
 5b,c).
-Lastly, MMD is not robust across different dimensions for a Gaussian kernel with a fixed bandwidth.
-We could make the MMD robust across dimensions by using median heuristic (Section 2.3), where we would increase the chosen bandwidth such that it stays on the order of the euclidean distance between datapoints.
-Note that MMD can in general be highly sensitive to its hyperparameters, and an appropriate value depends not only on the dimensionality of the data (Supp.
+Lastly, while here MMD is not robust across different dimensions for a Gaussian kernel with a fixed bandwidth, one typically chooses the bandwidth to be on the order of the distance between datapoints (e.g., using the median heuristic; Section 2.3).
+Note that in general, MMD can be highly sensitive to kernel- and hyperparameter choice; an appropriate setting depends not only on the dimensionality of the data (Supp.
 Fig.
-S5 and S6), but also on the structure of the distributions (Supp.
+S6, S7, and S8), but also on the structure of the distributions (Supp.
 Fig.
-S4).
+S5).
+SW distance, on the other hand, appeared robust to number of random projections used (Supp.
+Fig.
+S1 and A.4).
 3.2 FID-like distance comparison on ImageNet
+To explore scaling properties of FID-like distances, we generated high-quality synthetic samples using a state-of-the-art diffusion model as described by Dockhorn et al.
+(2022), which we compared to images from the ImageNet dataset, which contains 1000 classes, 100 images per class.
+All images were embedded using the pre-trained InceptionV3 network (Deng et al., 2009, Szegedy et al., 2016), following the implementation of Heusel et al.
+(2017), transforming the raw images into a 2048-dimensional feature space.
 Figure 9: Comparison of distances for ImageNet.
-(a) A comparison between the ImageNet test set and samples generated by an unconditional diffusion model, with varying sample sizes.
+(a) A comparison between the ImageNet test set and samples generated by an unconditional diffusion model with varying sample sizes.
 (b) Distance evaluation on dog classes (D) versus non-dog classes (~D), highlighting differences in image representation between these two categories of real data.
-(c) Distances between sets of randomly selected images vs varying number of included classes of images (from 10 to 1,000) from the test set, using synthetic samples created by a conditional diffusion model.
-To also explore scaling properties of FID-like distances, we used images from the ImageNet dataset and embedded them with the InceptionV3 network (Deng et al., 2009, Szegedy et al., 2015b), following the implementation of Heusel et al.
-(2017).
-In addition to the 100,000 images in the ImageNet test dataset (1000 classes, 100 images per class), we generated high-quality synthetic samples using a state-of-the-art diffusion model as described by Dockhorn et al.
-(2022).
-We first produced 50,000 samples with the base unconditional version of this model.
-Using a conditional generative model, we additionally generated 100,000 class-conditional images (i.e., 100 per class), exactly matching the class distribution of the test-set.
-All of these images were embedded using the pre-trained InceptionV3 network (Szegedy et al., 2015a), transforming the raw images into a 2048-dimensional feature space.
+(c) Distances between sets of randomly selected images vs.
+varying number of included classes of images (from 10 to 1,000) from the test set, using synthetic samples created by a conditional diffusion model.
+Sample size
+We first produced varying number of samples with the base unconditional version of the diffusion model and compared those to samples from Imagenet (Fig.
+9a).
 Calculating the FID involves computing the mean and covariance of the distributions in the embedding space and then calculating the squared Wasserstein distance analytically.
 However, we broadened our evaluation by applying additional distances to the distributions in the embedding space.
-While SW distance, C2ST, and FID effectively highlight the greater dissimilarity of synthetic samples to real images (i.e., the ImageNet test set) ( Fig.
-9a) even for low sample sizes, the distinctiveness of the FID becomes only apparent when analyzing more than around 2000 samples.
-Estimating the full covariance matrix of the 2048-dim features with fewer samples leads to degeneracy and, thus, to numerical issues computing the square root.
-Common implementations hence generally recommend using more than 2048 samples (Heusel et al., 2018a; b).
-As also shown by Jayasumana et al.
-(2023), Betzalel et al.
-(2022), the Gaussian assumption in the FID is violated and can lead to problematic behavior.
+While SW distance, C2ST, and FID effectively highlight the greater dissimilarity of synthetic samples to real images (i.e., the ImageNet test set) even for low sample sizes, the distinctiveness of the FID only becomes apparent when analyzing more than 2000 samples.
+This is in line with common implementations which generally recommend using at least 2048 samples (Heusel et al.
+2018a; b).
+Another reason for the limited performance could be that the Gaussian assumption in the FID can be violated (Jayasumana et al., 2023, Betzalel et al., 2022).
 In contrast, the other distances reliably estimate a larger inter-dataset distance in regimes with few samples.
-In our subsequent analysis, we aimed to determine the effectiveness of various distances in discriminating between images from different classes by comparing the distances between different levels in the WordNet hierarchy (Miller, 1995).
-To this end, we focused on comparing images of dogs (D) with those of non-dog (~D) images (Fig.
-9b).
+Class discriminability
+We aimed to determine the effectiveness of various distances in discriminating between images from different classes.
+To this end, we focused on comparing ImageNet images of dogs (D) with ImageNet images of other (non-dog) classes (~D; Huang et al.
+2021).
 All investigated distances, except FID, were successful in identifying images from different classes as being more distinct than images from the same class.
@@
 ~D).
-This shows that comparing two image classes can be problematic with FID, which is usually used to compare two distributions over many image classes (i.e.
-over natural images).
-To examine the effects only including a subset of classes from the dataset (i.e., modes of the distribution), we employed a test set and a conditional synthetic dataset, each comprising 1,000 classes with 100 samples per class.
-Our analysis involved comparing the complete test set against synthetic datasets that included only subsets of classes (Fig.
+This indicates that FID may be comparably insufficient in recognizing dissimilarities between distributions that have clear distinctions (whether across individual classes or in a one vs.
+all scenario).
+Mode coverage
+To examine the effects when only considering a subset of classes from the dataset (i.e., modes of the distribution), we used the diffusion model to generate class-conditional images matching the classses of the ImageNet test set (Fig.
 9c).
-For comparison purposes and as a control measure, we also conducted a scenario where, instead of selectively excluding classes, we randomly removed an equivalent number of images from the dataset.
-This approach revealed that limiting the dataset to a small number of classes compromised the performance across all evaluated distances, in contrast to the outcomes observed when randomly excluding a subset of images.
+Our analysis involved comparing the complete test set against model-generated datasets that included only subsets of classes.
+For comparison, we created a control dataset by randomly sampling from all classes (dashed curve in Fig.
+9c).
+This approach revealed that limiting the dataset to a small number of classes compromised the performance across all evaluated distances, in contrast to the outcomes observed when randomly excluding a subset of images—indicating that all distances successfully detected the lack of mode coverage.
 To achieve performance comparable to that observed with random removals, it was necessary to include at least 800 classes in the comparison.
-As the InceptionV3 network is, in essence, trained to classify ImageNet images (Szegedy et al., 2015a) (under certain regularization schemes), hence the extracted high-level features may also be very sensitive to class-dependent image features and not necessarily for general image quality.
+As the InceptionV3 network is, in essence, trained to classify ImageNet images (Szegedy et al.
+2015; under certain regularization schemes), the extracted high-level features may also be very sensitive to class-dependent image features and not necessarily for general image quality.
 This behavior can be observed in Fig.
@@
 (2023).
-By replacing InceptionV3 with other embedding networks (e.g.
-CLIP which is trained to match images to captions) this class sensitivity can be reduced (Kynkäänniemi et al., 2022).
-3.4 Computational complexity
-While we only considered the sensitivity and specificity of the different distances in the previous paragraphs, we want to highlight that they differ also in their computational complexity.
-With respect to number of samples $N$, MMD and FID have a complexity of $O(N^{2})$ (note that for MMD the computational cost can be reduced, potentially at the cost of making approximation; Gretton et al.
-2012a, Zhao & Meng 2015, Cheng & Xie 2021, Bodenham & Kawahara 2023, Bharti et al.
-2023, Gretton et al.
-2012b).
-While SW distance scales with $O(N\log N)$ (Nadjahi, 2021), it is difficult to make principled assessments of the computational complexity for C2ST, as it is highly dependent on the chosen classifier.
-But as more samples lead to larger training and test datasets, sample size is likely to influence the compute time.
-Similarly, the computational complexity of computing these distances increases as the dimensionality of the data increases, with non-trivial scaling depending on the task and hyperparameters chosen.
-Note that despite their differences, all presented distances are reasonably tractable in the settings of our experiments, whereas the scaling experiments might be computationally unfeasible for other distances or datasets.
-We therefore strongly recommend carefully considering the complexity of the measure before conducting experiments on high-dimensional or very large datasets.
+By replacing InceptionV3 with other embedding networks (e.g., CLIP, which is trained to match images to corresponding text), this class sensitivity could be reduced (Kynkäänniemi et al., 2023).
+Additional generative models
+We additionally generated images using a consistency model (CM) for unconditional image generation (Song et al., 2023), to investigate how these metrics perform when evaluated on images created by different generative models.
+This model was trained on ImageNet 64x64, similar to the GENIE model (Dockhorn et al., 2022).
+Moreover, we included the following models: BigGAN (Brock et al., 2019), ablated diffusion model (ADM; Dhariwal & Nichol 2021), Glide (Nichol et al., 2022), Vector Quantized Diffusion Model (VQDM; Gu et al.
+2022), Wukong (Wukong, 2022), Stable diffusion 1.5 (SD1.5; Rombach et al.
+2022b) and Midjourney (Midjourney 2022; details in Appendix A.13).
+We evaluated the metrics (and multiple commonly used variants of KID and C2ST) for each model against the ImageNet test set (see Table S4).
+While there is some agreement between the metrics regarding which model generates images closest to the ImageNet test set, there are also differences in the relative ordering across different metrics.
+As expected, the most recent unconditional models trained directly on ImagenNet 64x64 performed best in our evaluation (GENIE, CS), better than the two other unconditional generative models (BigGAN, ADM).
+The other models are text-to-image and thus only prompted to generate images from specific ImageNet classes (GLIDE, VQDM, Wukong, SD1.5, Midjourney).
+Interestingly, the prompted models performed better than older unconditional models (BigGAN, ADM) most of the time.
+Recall that all we evaluate is the similarity to the ImageNet test set; prompted versions might produce images from the correct classes but might contain differences in style or appearance compared to actual images in ImageNet.
+Despite the demonstrated class-sensitivity (Fig.
+9c), the InceptionV3 embeddings are thus also sensitive to different “styles” of natural images.
+Note that in this case, being closer to ImageNet does not necessarily mean generating better images (based on human perception), but rather creating images that are more ImageNet-like.
 4 Scientific applications
@@
 For each application, we used two generative models or simulators to sample synthetic data.
-We then compared the synthetic samples to real data using the discussed distances.
+We then compared the synthetic samples to real data (hold-out test set) using the discussed distances.
 To obtain baseline values for each distance, we computed distances between subsets of real data.
@@
 (a) Schematic of a Drift-Diffusion model (DDM), a classical neuroscientific model of decision making behavior.
-Overall evidence drives the model toward one of two choices (drift), but sensory and environmental noise result in random fluctuations in evidence integration (diffusion).
-(b) Distributions of primate decision times from the real dataset (black), and two fitted models of varying complexity: DDM1 (gray) and DDM2 (gold).
-(c) SW distance, C2ST, MMD (bandwidth=0.5) between subsets of the three distributions.
+Overall, evidence drives the model toward one of two choices (drift), but sensory and environmental noise result in random fluctuations in evidence integration (diffusion).
+(b) Distributions of primate decision times from the test set (black), and two fitted models of varying complexity: DDM1 (gray) and DDM2 (gold).
+(c) SW distance, C2ST, MMD (bandwidth=0.5) between subsets of the generated and real data distributions.
 FID is not applicable in these comparisons, because the data are one-dimensional distributions.
@@
 We evaluated two versions of a Drift-Diffusion Model (DDM; Fig.
-10a) (Ratcliff, 1978), a frequently used model in cognitive neuroscience.
+10a;  Ratcliff 1978), a frequently used model in cognitive neuroscience.
 The two versions differ with respect to the drift rate, which is the speed and direction at which evidence accumulates towards a decision, and the decision boundaries, which determine how much evidence is needed to make a decision.
-Specifically the first version (DDM1) uses a drift rate that varies linearly with position and time, and decision boundaries that decay exponentially over time whereas the second version (DDM2) uses a drift rate and decision boundaries that are constant over time (for details, see Section A.9).
-We fitted each model against empirical primate decision times with the use of the pyDDM toolbox (Shinn et al., 2020), generated one-dimensional synthetic datasets, and compared each dataset to the actual primate decision time distributions.
+Specifically, the first version (DDM1) uses a drift rate that varies linearly with position and time and decision boundaries that decay exponentially over time, whereas the second version (DDM2) uses a drift rate and decision boundaries that are constant over time (for details, see Section A.14).
+We fitted each model to empirical primate decision times with the use of the pyDDM toolbox (Shinn et al., 2020), generated one-dimensional synthetic datasets, and compared each dataset to the actual primate decision time distributions.
 While the resulting distributions of decision times are visually similar (Fig.
@@
 10c).
-For C2ST DDM1 and the real data distribution are even indistinguishable, with median C2ST values around $0.5$.
+For C2ST, DDM1 and the real data distribution are even indistinguishable, with median C2ST values around $0.5$.
 This suggests that SW and MMD provide a more nuanced differentiation between the models.
@@
 Figure 11: Comparing generated and real X-ray images.
-(a) Sketch of the embedded distributions of X-ray images from the three different datasets: real dataset (black), Progressive Growing Generative Adversarial Network (PGGAN) (grey), and Stable Diffusion (SD) model (gold).
+(a) Sketch of the embedded distributions of X-ray images from the three different datasets: test set of real dataset (black), Progressive Growing Generative Adversarial Network (PGGAN) (grey), and Stable Diffusion (SD) model (gold).
 (b) Examples from real and generated X-ray images.
 Three full-view examples from each distribution and four examples magnifying the top right corner.
-(c) SW distance, C2ST, MMD (bandwidth=50), and FID between samples of the three distributions of embedded X-ray images.
+(c) SW distance, C2ST, MMD (bandwidth=50), and FID between samples of the real and generated distributions of embedded X-ray images.
 Scatter-points indicate comparisons between ten random subsets from each dataset.
 Thick horizontal bars indicate median values.
-In the second application we turned to a high dimensional example, in which we compared synthetic X-ray images generated by a Progressive Growing Generative Adversarial Network (PGGAN) model  (Segal et al., 2021) and by a StableDiffusion (SD) model (Malik & Humair, 2023) to real chest X-ray images from the ChestX-ray14 dataset (Wang et al., 2017).
+In the second application we turned to a high-dimensional example, in which we compared synthetic X-ray images generated by a Progressive Growing Generative Adversarial Network (PGGAN) model (Segal et al., 2021) and by a StableDiffusion (SD) model (Malik & Humair, 2023) to real chest X-ray images from the ChestX-ray14 dataset (Wang et al., 2017).
 Each image has a total dimension of $1024\times 1024$ pixels.
@@
 11b).
-For instance, in real images, the top often contains annotations including e.g.
-patient id, side of the body, or the date the X-ray was taken.
+In real images, the top often contains annotations including, e.g., patient id, side of the body, or the date the X-ray was taken.
 These textual elements often contain artifacts or, in case of SD, are completely unrealistic.
 To compare these high-dimensional images, we embedded them in a 512-dimensional embedding space using the CheXzero network (Tiu et al., 2022), a CLIP (Radford et al., 2021) network fine-tuned for chest X-ray images.
-We opted for using this specialized network instead of the standard InceptionV3 network as it might overcome biases introduced by classification task training (Kynkäänniemi et al., 2022).
+We opted for using this specialized network instead of the standard InceptionV3 network as it might overcome biases introduced by classification task training (Kynkäänniemi et al., 2023).
 As expected, samples generated by PGGAN are closer to the real data across all distances compared to SD-generated data (Fig.
-11c), likely due to the unrealistic sharpness and more obvious textual artifacts of the SD-generated images.
-However, C2ST is even high between PGGAN outputs and the real data, suggesting that the high-dimensionality of the data increases the sensitivity of this measure.
+11c), likely due to unrealistic sharpness and more obvious textual artifacts of the SD-generated images.However, C2ST is even high between PGGAN outputs and the real data, suggesting that the high-dimensionality of the data increases the sensitivity of this measure.
 Taken together, our results suggest that PGGAN is more accurate in generating realistic X-ray images compared to SD.
-Our findings highlight that using different metrics can support different conclusions.
+Our findings show that using different metrics can support different conclusions.
 For instance, C2ST suggests equality between DDM1 and real decision time data, whereas SW distance and MMD metrics indicate a larger difference between DDM1 and the real data.
@@
 5 Discussion
-This work describes and explores four commonly applied sample-based distances representing different methodologies for defining statistical distance: Using low-dimensional projections (SW), obtaining a distance using classifiers (C2ST), using embeddings through kernels (MMD) or neural networks (FID).
+This work describes and characterizes four commonly applied sample-based distances representing different methodologies for defining statistical distance: Using low-dimensional projections (SW), obtaining a distance using classifiers (C2ST), using embeddings through kernels (MMD) or neural networks (FID).
 Despite their operational differences, they are all based on a fundamental concept: simplifying complex distributions into more manageable feature representations to facilitate comparison.
@@
 There are approaches to reduce this effect by considering other projections than simple linear slices, as described in Section 2.1.
-Due to its computational efficiency and differentiability, the SW distance can also be used as a loss function to train generative models (Wu et al., 2019, Deshpande et al., 2018; 2019, Liutkus et al., 2019, Vetter et al., 2024).
-In our experiments, the metric did show convincing results and in contrast to the MMD, C2ST, and FID, SW distance does not require to choose specific hyperparameters for which results can differ drastically.
-Although currently not extensively used in literature for evaluation, this makes the SW distance efficient, scalable, and an objective baseline for general distribution comparisons.
-Yet, this also makes it less flexible to adapt to specific features of interest.
-Although, the majority of research on sliced distances focus on sliced Wasserstein metrics, slicing other metrics is also possible.
-For a certain subset of choices, equivalence to MMDs can be established (Kolouri et al., 2019).
+In our experiments, the metric did show convincing results and in contrast to the MMD, C2ST, and FID, SW distance does not require one to choose specific hyperparameters (for which results can differ drastically).
+Although currently not extensively used in literature for evaluation, this makes the SW distance efficient, scalable, and a consistent baseline for general distribution comparisons.Yet, this also makes it less flexible to adapt to specific features of interest.
+The Wasserstein and SW distances are not interpretable, and admit only biased sample-based estimates.
+This can be a limitation for some tasks.
+However, due to its computational efficiency and differentiability, the SW distance is commonly used as a loss function to train generative models, such as GANs (Deshpande et al., 2018, Wu et al., 2019), Autoencoders (Wu et al., 2019), nonparametric flows (Liutkus et al., 2019), normalizing flows (Dai & Seljak, 2021), and multi-layer perceptrons (Vetter et al., 2024).
+Although the majority of research on sliced distances focuses on sliced Wasserstein metrics, slicing other metrics is also possible.
+For a certain subset of choices, equivalence to MMDs can be established (Feydy et al., 2019, Kolouri et al., 2019, Hertrich et al., 2024).
 Classifier Two-Sample Test (C2ST)
@@
 The effectiveness of MMD largely depends on the kernel function chosen (implicitly representing the feature space), which affects how well it can spot differences between various types of data.
-Inappropriate kernel choice can leave the metric insensitive to subtle differences in the distribution (Gretton et al., 2012b, Sriperumbudur et al., 2009) (see Fig.
+Inappropriate kernel choice can leave the metric insensitive to subtle differences in the distribution (Gretton et al.
+2012b, Sriperumbudur et al.
+2009; see Fig.
 8).
-The MMD can be estimated efficiently and is differentiable, and thus often used as a loss function for training generative models (Arbel et al., 2019, Li et al., 2017, Bińkowski et al., 2021, Briol et al., 2019).
+The MMD can be estimated efficiently and is differentiable, and thus often used as a loss function for training generative models (Dziugaite et al., 2015, Arbel et al., 2019, Li et al., 2017, Bińkowski et al., 2018, Briol et al., 2019).
 Yet, a kernel must satisfy certain criteria, e.g., positive definiteness, making the design of new kernel functions challenging.
@@
 (2022a) proposed a method to estimate MMD via a witness function that determines MMD (Appendix A.2).
-This method is closely related to C2ST in that both two estimate a discrepancy among distributions via a classifier (Kübler et al., 2022a, Section 5).
+This method is closely related to C2ST in that both estimate a discrepancy among distributions via a classifier (Kübler et al., 2022a, Section 5).
 Network-based
@@
 Even more important than for the C2ST, this network must be well-established and shared which is a necessary but not sufficient criterion (Chong & Forsyth, 2020) to compare different results.
-While such well-established defaults exist for images (Szegedy et al., 2015a, Radford et al., 2021), this is not the case for other domains.
-For example, the time series generation community did not yet establish a default, and embedding nets are either trained or chosen by the authors (Smith & Smith, 2020, Jeha et al., 2021).
-We showed that the class-sensitivity of FID (Section 3.3) tends to model collapse (such as GANs), but might not necessarily reflect general image quality.
-In fact, Betzalel et al.
+While such well-established defaults exist for images (Szegedy et al., 2015, Radford et al., 2021), this is not the case for other domains.
+For example, the time series generation community did not yet establish a default, and embedding networks are either trained or chosen by the authors (Smith & Smith, 2020, Jeha et al., 2021).
+We demonstrated that the class-sensitivity of FID (Section 3.2) often leads to model collapse, as seen in GANs.
+However, it may not accurately reflect the overall image quality.
+For example, Betzalel et al.
 (2022), Kynkäänniemi et al.
@@
 Yet, FID features have been shown to align much better with human perception than traditional metrics (Zhang et al., 2018).
+Recent developments in network-based approaches include the use of Central Kernel Alignment (CKA; Cortes et al.
+2012) to compute the distance between network-embedded samples.
+CKA scores show considerable stability when evaluated with different choices of network architectures and layers (Yang et al., 2023).
+Another newly introduced metric, Mauve (Pillutla et al., 2021, Liu et al., 2021, Pillutla et al., 2023), can, for instance, be used to measure how close machine-generated text is to human language using an external language model to embed the samples from each distribution.
+This metric uses divergence frontiers to take into account the trade-off between quality and diversity when evaluating generative models.
 Closing remarks
-Ultimately, the choice of distance hinges on the nature of the data under consideration and the specific characteristics of it one aims to compare.
-Given a specific dataset and problem, one will likely have to look beyond the distances discussed in this paper.
-For example, in the realm of human-centric data like images and audio, the perceptual indistinguishability of distributions is important (Gerhard et al., 2013, Zhang et al., 2018).
-Time series data, characterized by its temporal structure, demand metrics that account for temporal shifts and variations in a manner that does not disproportionately penalize minor discrepancies in timing, such as Dynamic Time Warping (Müller, 2007), or by using frequency information (Hess et al., 2023).
-In general however, irrespective of the specific use-case, it is advisable to use multiple distances in order to obtain a full picture, as using a single distances individually could support competing conclusions about the model that is to be evaluated.
-Throughout this paper we have explained and analyzed four approaches of measuring statistical distance.
-While this is only a small subset of all possible distances available, we hope to have provided the foundational knowledge with which researchers can find, understand and interpret statistical distances specific to their own scientific application.
+Ultimately, the choice of distance hinges on the nature of the data under consideration and the specific characteristics one aims to compare.
+Given a particular dataset and problem, it may be necessary to look beyond the distances discussed in this paper.
+For instance, in the realm of human-centric data such as images and audio, perceptual indistinguishability of distributions is crucial (Gerhard et al., 2013, Zhang et al., 2018).
+Time series data, with its inherent temporal structure, requires metrics that accommodate temporal shifts and variations without disproportionately penalizing minor discrepancies in timing, such as Dynamic Time Warping (Müller, 2007) or frequency-based methods (Hess et al., 2023).
+Additional recent works propose new distances based on fields such as topology (Barannikov et al.
+2021).
+In general, regardless of the specific use case, it is advisable to use multiple distance measures to obtain a comprehensive view, as relying on a single measure may lead to competing conclusions about the model under evaluation.
+Throughout this paper, we have explained and analyzed four approaches to measuring statistical distance.
+While this represents only a small subset of all possible measures, we hope to have provided the foundational knowledge researchers need to find, understand, and interpret statistical distances relevant to their own scientific applications.
 Code availability
@@
 Writing—Original Draft: AS, CS, FG, FP, GM, JK, JV, LH, MD, MG, MP, RG, RR, SB, ZS
-Writing — Review & Editing: AD, AS, CS, FP, GM, JHM, JKL, JV, LH, LU, MG, MP, SB, ST, ZS
+Writing—Review & Editing: AD, AS, CS, FP, GM, JHM, JKL, JV, LH, LU, MG, MP, RG, SB, ST, ZS
 Visualization: AD, AES, AS, FP, GM, JKL, JV, MD, MG, MP, RG, RR, ZS
@@
 This work was supported by the German Research Foundation (DFG) through Germany’s Excellence Strategy (EXC-Number 2064/1, PN 390727645) and SFB1233 (PN 276693517), SFB 1089 (PN 227953431), SPP2041 (PN 34721065), the German Federal Ministry of Education and Research (Tübingen AI Center, FKZ: 01IS18039; DeepHumanVision, FKZ: 031L0197B; Simalesam: FKZ 01|S21055A-B), the Carl Zeiss Foundation (Certification and Foundations of Safe Machine Learning Systems in Healthcare), the Else Kröner Fresenius Stiftung (Project ClinbrAIn), and the European Union (ERC, DeepCoMechTome, 101089288).
-ST was partially supported by Grant-in-Aid for Research Activity Start-up 23K19966 and JST CREST JPMJCR2015.
-The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting SB, MD, MG, JK, JL, GM, MP, RR, AS, ZS and JV.
+ST was partially supported by Grant-in-Aid for Research Activity Start-up 23K19966, Grant-in-Aid for Early-Career Scientists 24K20750 and JST CREST JPMJCR2015.
+SB, MD, MG, JK, JL, GM, MP, RR, AS, ZS and JV are members of the International Max Planck Research School for Intelligent Systems (IMPRS-IS).
 Appendix A Appendix
 A.1 Generative Models in Science
-Table S1: Example generative models in science.
-Simulators include mechanistic models, DNN based models e.g.
-VAEs, GANs, Diffusion models.
-This is not an exhaustive list regarding disciplines using generative models nor generative models used in the listed disciplines.
-ML model Simulator
-natural science biology
+Table S1: Example generative models in science. This is not an exhaustive list regarding disciplines using generative models nor generative models used in the listed disciplines.
+Discipline Model
+Biology
+[…]
 […]
@@
 where $p_{1}(x)$ and $p_{2}(y)$ are the probability distributions of random variables $x,y\in\mathcal{X}$, and $\phi:\mathcal{X}\rightarrow\mathcal{H}$.
-Generally $\mathcal{X}$ and $\mathcal{H}$ are defined as a topological space, and the reproducing kernel Hilbert space (RKHS), respectively, but readers can simply think of the euclidean space $\mathbb{R}^{N}$ for the first examples in the main text.
+Generally, $\mathcal{X}$ and $\mathcal{H}$ are defined as a topological space, and the reproducing kernel Hilbert space (RKHS), respectively, but readers can simply think of the Euclidean space $\mathbb{R}^{N}$ for the first examples in the main text.
 For the identity feature map $\phi^{(1)}:\mathbb{R}\rightarrow\mathbb{R},\phi^{(1)}(x)=x$, MMD can be computed as
@@
 For given distribution $p(x)$, the kernel mean embedding $\mathbb{E}_{p(x)}[k(x,u)]\in\mathcal{H}$ is an element in RKHS that satisfies $\langle\mathbb{E}_{p(x)}[k(x,u)],f(u)\rangle_{\mathcal{H}}=\mathbb{E}_{p(x)}[f(x)]$ for any $f\in\mathcal{H}$ with argument $u\in\mathcal{X}$.
-The embedding $\mathbb{E}_{p(x)}[k(x,u)]$ is known to be determined uniquely if a corresponding kernel is bounded, i.e.
-$\|k(x,x^{\prime})\|_{\mathcal{H}}<\infty$ for any $x$.
+The embedding $\mathbb{E}_{p(x)}[k(x,u)]$ is known to be determined uniquely if a corresponding kernel is bounded, i.e., $\|k(x,x^{\prime})\|_{\mathcal{H}}<\infty$ for any $x$.
 Then, as shown in Gretton et al.
@@
 Note that linear and polynomial kernels are not characteristic, while they are quite popular in natural language processing.
-A.3 Details about Wasserstein and Sliced-Wasserstein distances
-Formally, the Wasserstein distance is described in measure-theoretic terms.
-We first state this definition, and then provide an accessible interpretation in the common case that the measures have well-defined probability density functions.
+A.3 Formal Definition of Wasserstein and Sliced-Wasserstein Distance
 Let $(M,\rho)$ be a Polish space, $\mu,\nu\in P(M)$ be two probability measures, and let $q\in[1,+\infty)$. Then the Wasserstein-$q$ distance between $\mu$ and $\nu$ is defined as
@@
 where $\Gamma(\mu,\nu)$ is the set of all couplings between $\mu,\nu$.
-Consider the case where $M\subseteq\mathbb{R}^{d}$, $\rho$ is the $q$-norm $||\cdot||_{q}$, and $\mu$ and $\nu$ have well-defined probability density functions $p_{1}$ and $p_{2}$ respectively.
-Then the Wasserstein-$q$ norm can be written as
-$W_{q}(p_{1},p_{2})=\inf_{\gamma\sim\Gamma(p_{1},p_{2})}\left(\mathbb{E}_{x_{1},x_{2}\sim\gamma}\,||x_{1}-x_{2}||_{q}^{q}\right)^{\frac{1}{q}},$
-where $\Gamma(p_{1},p_{2})$ is the set of all couplings, that is all possible “transportation plans”, between $p_{1}$, $p_{2}$.
-$\gamma\in\Gamma(p_{1},p_{2})$ is a joint distribution over $(x_{1},x_{2})$ with respective marginals $p_{1}$ and $p_{2}$ over $x_{1}$ and $x_{2}$.
-The Sliced-Wasserstein distance is similarly defined in measure-theoretic terms for the measures $\mu$ and $\nu$.
-We refer the reader to Nadjahi (2021) for details.
-In the less general case described above, we can similarly provide a more intuitive definition.
-In particular, the random projection directions described in 2.1 are uniformly random vectors $u\in\mathbb{S}^{d-1}$, the unit sphere in $\mathbb{R}^{d}$.
-Projecting the distributions $p_{1}$ and $p_{2}$ onto $u$ induces one-dimensional distributions $p_{1}^{u}$ and $p_{2}^{u}$ with samples $u^{\mathsf{T}}x_{i}$, where $x_{1}\sim p_{1}$ and $x_{2}\sim p_{2}$.
+The Sliced-Wasserstein distance is closely connected to the Radon transform (Helgason, 1980).
+We direct readers to Bonneel et al.
+(2015) for details.
+Here we provide a shorter definition following Definition 2.9 of Nadjahi (2021).
+Suppose that $M\subset\mathbb{R}^{d}$, and denote by $\mathbb{S}^{d-1}=\{\theta\in\mathbb{R}^{d}\,:\,||\theta||_{2}=1\}$ be the unit sphere with respect to the Euclidean norm.
+Let $u^{*}:X\to\mathbb{R}$ be a the linear form given by $u^{*}(x)=\langle u,x\rangle$, and $q\in[1,+\infty)$.
+Then the Sliced-Wasserstein distance (of order $q$) is defined for any measures $\mu,\nu\in P_{q}(X)$ as
+[…]
+Where $\mathcal{U}$ is the uniform distribution on $\mathbb{S}^{d-1}$, and for any $u\in\mathbb{S}^{d-1}$, $u_{\#}^{*}$ denotes the push-forward operator of $u^{*}$.
+As for the Wasserstein distance, the definition becomes more intuitive in the case where $\mu$ and $\nu$ admit the probability density functions ($p_{1}$ and $p_{2}$ respectively).
+In particular, the random projection $u$ are uniformly random vectors in $\mathbb{S}^{d-1}$.
+Therefore, projecting the distributions $p_{1}$ and $p_{2}$ onto $u$ induces one-dimensional distributions $p_{1}^{u}$ and $p_{2}^{u}$ with samples $u^{\mathsf{T}}x_{i}$, where $x_{1}\sim p_{1}$ and $x_{2}\sim p_{2}$.
 The Sliced-Wasserstein-$q$ distance can then be written as
-$SW_{q}(p_{1},p_{2})=\mathbb{E}_{u\sim\mathcal{U}(\mathbb{S}^{d-1})}[W_{q}(p_{1}^{u},p_{2}^{u})],$
-where $\mathcal{U}(\mathbb{S}^{d-1})$ is the uniform distribution over vectors on the unit sphere $\mathbb{S}^{d-1}$.
-A.4 C2ST scores below 0.5
+$SW_{q}(p_{1},p_{2})=\left(\mathbb{E}_{u\sim\mathcal{U}(\mathbb{S}^{d-1})}[W_{q}^{q}(p_{1}^{u},p_{2}^{u})]\right)^{1/q}.$ (7)
+A.4 Dependence of SW Distance on Number of Projections
+Figure S1: The SW distance estimate is not strongly sensitive to the number of random projections.
+We compare the SW distance estimate for the $\{10,100,1000\}$-dimensional Gaussian task with 1 shifted dimension (Sec.
+3) as we increase the number of random projections used in the estimation.
+As the number of projections increases, the variance of the SW distance estimate decreases, but across all dimensionalities considered, the SW distance estimate has converged by 100 random projections.
+All SW distance experiments in the main text were performed with the SW distance approximate with 100 random projections to approximate the expectation Eq.
+(7).
+Here, we additionally show the dependence of the SW distance approximation with finite projections on the $d$-dim Gaussian example (see Sec.±3), for $d\in\{10,100,1000\}$ (Fig S1).
+While the sample-based approximation to the analytic 1-dimensional Wasserstein distance is biased (Sec.
+2.1), the Monte Carlo approximation to the expectation Eq.
+(7) is an unbiased estimate of the sample-based Wasserstein distance.
+We observe that for very high (1000)-dimensional distributions, the SW distance estimate converges quickly, and the choice of 100 random projections for the computation of the SW distance is appropriate.
+A.5 C2ST scores below 0.5
 In practice the C2ST score can sometimes turn out to be below .5.
@@
 This effect is particularly noticeable if the classifier was not carefully regularized during training, and thus memorized the class label of the duplicate in the training set.
-A.5 Sliced-Wasserstein, MMD and C2ST as optimization target
+A.6 Sample and computational complexity
+Table S2: Summary of computational complexity and theoretical properties of metrics in terms of number of samples $N$ and data dimensionality $D$.
+Sample complexity here refers to the convergence rate of the sample-based estimate to the true value of the metric.
+*Bound based on Ghosal & Sen (2019), Nguyen & Ho (2024) for SWD and Gretton et al.
+(2012a) for MMD.
+Best case scenario.
+In practice, the computational cost of training a neural network scales superlinearly with both sample size and data dimensionality.
+General case (see Sec.
+2.3, A.6 for details).Cost of calculating the square root of the covariance matrix in Eq.
+4.
+SW C2ST MMD FID
+Sample Complexity ($N$) $\mathcal{O}(N^{-1/2})$* N/A $\mathcal{O}(N^{-1/2})$* N/A
+Computational Complexity ($N$) $\mathcal{O}(N\log N)$ $\mathcal{O}(N)^{\dagger}$ $O(N^{2})^{\ddagger}$ $\mathcal{O}(N^{2})$
+Computational Complexity ($D$) $\mathcal{O}(D)$ $\mathcal{O}(D)^{\dagger}$ $\mathcal{O}(D)$ $\mathcal{O}(D^{3})$
+Estimator unbiased? no no yes no
+We here highlight that the different distances also differ in their sample and computational complexities (Table S2).
+With respect to the number of samples $N$, MMD and FID have a complexity of $O(N^{2})$.
+Note that for MMD the computational cost can be reduced, e.g., by increasing variance or for a specific kernel choice (Gretton et al., 2012a, Zhao & Meng, 2015, Cheng & Xie, 2021, Bodenham & Kawahara, 2023, Bharti et al., 2023, Gretton et al., 2012b)).
+While SW distance scales with $O(N\log N)$ (Nadjahi, 2021), it is difficult to make principled assessments of the computational complexity for C2ST, as it is highly dependent on the chosen classifier.
+However, as more samples lead to larger training and test datasets, the sample size is likely to influence the compute time.
+Similarly, the computational complexity of computing these distances increases as the dimensionality of the data increases, with non-trivial scaling depending on the task and chosen hyperparameter.
+We also report the theoretical convergence of sample-based estimates for MMD, as well as SW, which is subject to active research.
+We report bounds from recent works (Ghosal & Sen, 2019, Gretton et al., 2012a).
+We do not report sample complexities for C2ST and FID, as these strongly depend on the choice of classifier and embedding network, respectively.
+Note that despite their differences, all presented distances are reasonably tractable in the settings of our experiments, whereas the scaling experiments might be computationally unfeasible for other distances or datasets.
+Therefore, we strongly recommend carefully considering the complexity of the measure before conducting experiments on high-dimensional or very large datasets.
+We report the practical computation times for our experiments in Appendix A.11.
+A.7 Mode coverage properties
+Mode coverage is the ability of a model to capture and generate diverse data, i.e., from multiple modes of the underlying distribution (see Fig.
+S2;  S3).
+The community has focused on evaluation of mode coverage with different metrics driven by the development of GANs (Goodfellow et al., 2014, Gui et al., 2023, Saad et al., 2024).
+Empirically, in training generative models, mode coverage has been found to trade off with sample quality and speed, illustrating the generative learning trilemma (Xiao et al., 2022).
+All presented metrics can distinguish between a collapsed and a full distribution (Fig.
+9), an empirical finding also reported in previous work (Che et al., 2017, Li et al., 2017, Deshpande et al., 2018, Borji, 2019).
+However, their sensitivity relies on different factors: SWD captures different modes when they are separated in one-dimensional projections, MMD depends on appropriate kernel choice, FID on expressive embeddings, and C2ST on well-trained classifiers.
+Other metrics used to quantify mode collapse include precision and recall (Kynkäänniemi et al., 2019).
+A.8 Sliced-Wasserstein, MMD and C2ST as optimization target
 Figure S2: Trade-offs illustrated through optimization of a miss-specified model.
@@
 The models optimised with C2ST and MMD are mode-seeking: they have high densities only in the largest mode of the true distribution, and thus produces likely, but unvaried samples.
-A.5.1 Fitting a Gaussian with gradient descent
+A.8.1 Fitting a Gaussian with gradient descent
 We provide an illustrative example of which distributions are obtained when using Wasserstein, MMD and C2ST distances as a goodness of fit criterion, for both a miss-specified example, in Fig.
-S1 and well-specified example Fig.
-S2.
+S2 and well-specified example Fig.
+S3.
 We can see that in the miss-specified example different distances make different trade-offs, for example whether they are mode-seeking, and produce likely but unvaried samples, or are mode-covering, where they produce varied, but also potentially unlikely samples.
@@
 We used the ADAM optimizer (Kingma & Ba, 2015), with learning rate=0.01 and default momentums, using 2500 epochs of 10000 samples.
-A.5.2 Fitting a mixture of Gaussians with Expectation-Maximisation
+A.8.2 Fitting a mixture of Gaussians with Expectation-Maximisation
 We also include an example where the model we fit is well-specified, which in this case means it is also a mixture of two Gaussians (Fig.
-S2.
+S3.
 As directly optimizing a mixture distribution with gradient descent is not straightforward, we used an Expectation-Maximization algorithm (where we use the distances instead of the log-likelihood in the maximisation step)
@@
 We then iteratively performed the following two steps to optimise the model.
-E-step: For each of the $N$ datapoints $\hat{x}_{i}$ from $p_{\text{true}}$, we calculated the probability of it belonging mixture component 1 or 2:
+E-step: For each of the $N$ datapoints $\hat{x}_{i}$ from $p_{\text{true}}$, we calculated the probability of it belonging to mixture component 1 or 2:
 $p(z=k|\hat{x}_{i})=\frac{p(\hat{x}_{i}|z=k)p(z=k)}{\sum_{k^{\prime}}^{2}p(\hat{x}_{i}|z=k^{\prime})p(z=k^{\prime})}.$
@@
 Plotted are the contour lines of .25, .75, 1, and 2.5 standard deviations of the fitted Gaussians multiplied by their corresponding mixture weight.
-A.6 Additional scaling experiments with different sample size budgets and ranges
+A.9 Additional scaling experiments with different sample size budgets and ranges
 In Fig.
@@
 Here, we further examine the performance of the distances across datasets of varying sample sizes, particularly for small sample set sizes, ranging from only 8 to 80 samples per set (Fig.
-S3).
-We examine three distinct data configurations where the distinction between the true and approximated distributions progressively increases from subpanels S3 a to c.
+S4).
+We examine three distinct data configurations where the distinction between the true and approximated distributions progressively increases from subpanels S4 a to c.
 Across all distances, it becomes evident that the larger the disparity between the two distributions, the fewer samples are needed for differentiation.
@@
 Fig.
-S3 a, b), all distances exhibit poor performance in distinguishing between the distributions.
+S4 a, b), all distances exhibit poor performance in distinguishing between the distributions.
 Figure S4: The larger the difference between two distributions the fewer samples suffice to tell the true and shifted distribution apart.
@@
 (c) Distances for a ten-dimensional standard normal distribution, for which all dimensions are shifted by one for the shifted example.
-A.7 Additional scaling experiments for different dimensionality of the data
+A.10 Additional scaling experiments for different dimensionality of the data
 When comparing the robustness of the measures with respect to the dimensionality of the data in Fig.
@@
 Extending this analysis, Fig.
-S4 presents similar experiments conducted on datasets where we compare an n-dimensional standard normal distribution with one where either all dimensions are mean-shifted by one (thus aligning with the C2ST experiment in Fig.
+S5 presents similar experiments conducted on datasets where we compare an n-dimensional standard normal distribution with one where either all dimensions are mean-shifted by one (thus aligning with the C2ST experiment in Fig.
 5b) or where all variances are increased by one.
 Fig.
-S4a corresponds to the experiment outlined in Fig.
+S5a corresponds to the experiment outlined in Fig.
 8c on dimensionality.
@@
 Figure S5: The impact of dimensionality varies across distances, with certain distances facing particular challenges in higher dimensional spaces.
-We compare sample sets of varying dimensionality (between 5 and 1000) of a ’true’ distribution either with a second dataset of the same distribution or with a sample set from an approximated/shifted distribution.
+We compare sample sets of varying dimensionality (between 5 and 1000) of a ‘true’ distribution either with a second dataset of the same distribution or with a sample set from an approximated/shifted distribution.
 The sample size is fixed to $10$k for all experiments and we show the mean and standard deviation over five runs of randomly sampled data.
@@
 (c) Distances for a sample set from an n-dimensional standard normal distribution, for which variances are increased by one for all dimensions.
-A.8 Sensitivity of the MMD bandwidth parameter
+A.11 Comparisons of practical compute times of different measures
+Before computing such measures, in particular for scaling experiments such as the ones presented here, where measures are calculated across a large range of sample sizes $N$ and data dimensions $D$, the practical compute time of the chosen measures should be considered.
+Depending on the downstream application, it might be time-critical to quickly evaluate distances which might favor some measures over others.
+Aligned with theoretical considerations regarding sampling complexity etc.
+as presented in Table S2), empirical compute times vary between the different measures.
+Given that empirical computational times for a single measure itself vary depending on the exact implementation, compute infrastructure, and problem at hand, we list approximated compute times for running the scaling experiments in Table S3.
+The calculated runtime combines both the comparisons of the ‘true‘ and the ‘shifted‘ or ‘approximated‘ experiments.
+Each experiment contains five repeats across different sampled data subsets.
+The sample size experiment contains eight different sample size values $N$ (50, 100, 200, 500, 1000, 2000, 3000, 4000), and the dimensionality experiment scales tests six different $D$ (5, 10, 50, 100, 500, 1000).
+For more details, see Section 3.
+The version of C2ST we use here, which is based on NN-based classifiers, takes orders of magnitude longer to compute than SW and MMD.
+Note, however, that alternative implementations and classifier variants could speed this up.
+Table S3: CPU wallclock run times for the comparison scaling experiments in Fig.8.
+The runtime combines both the comparisons of the true vs.
+the shifted or approximated distribution.
+Values are rounded estimates.
+sample size experiment dimensionality experiment
+2-dim MoGs 10-dim Gaussian n-dim Gaussian
+SWD 0.3 s 0.3 s 1.5 s
+C2ST 150 s 300 s 1500 s
+MMD 3 s 3 s 80 s
+A.12 Sensitivity of the MMD kernel choices and hyperparameters
+The general formulation of MMD allows for a wide range of kernel choices, each potentially with their own hyperparameters.
+These choices can have significant effect on its behavior.
+We provide some experiments demonstrating of the importance of well-tuned bandwidth parameters for Gaussian Kernel MMD as well as the impact of different kernels.
 Figure S6: The bandwidth parameter in Gaussian Kernel MMD is a sensitive parameter that requires careful selection for each dataset.
 The sample size is fixed to $10$k for all experiments and we show the mean and standard deviation over five runs of randomly sampled data.
-(a) $MMD^{2}$ distance with varying bandwidth parameters between 0.1 and 5 for the 2d-MoG example compared to samples from a unimodal Gaussian approximation with the same mean and covariance.
-(b) $MMD^{2}$ distance with varying bandwidth parameters between 0.1 and 20 for a 10-dimensional standard normal distribution, for which the first dimension is shifted by one for the shifted example.
-(c) $MMD^{2}$ distance with varying bandwidth parameters between 1 and 40 for a 100-dimensional standard normal distribution, for which the first dimension is shifted by one for the shifted example.
+(a) $\mathsf{MMD}^{2}$ distance with varying bandwidth parameters between 0.1 and 5 for the 2d-MoG example compared to samples from a unimodal Gaussian approximation with the same mean and covariance.
+(b) $\mathsf{MMD}^{2}$ distance with varying bandwidth parameters between 0.1 and 20 for a 10-dimensional standard normal distribution, for which the first dimension is shifted by one for the shifted example.
+(c) $\mathsf{MMD}^{2}$ distance with varying bandwidth parameters between 1 and 40 for a 100-dimensional standard normal distribution, for which the first dimension is shifted by one for the shifted example.
+We first vary the bandwidth parameter with fixed sample sizes for the three example datasets used in Section 3 (Fig.
+S6).
+We show that the estimated $\mathsf{MMD}$ values vary significantly across bandwidths, and both setting the bandwidth too low or too high yield poor results.
+However, we note that the values yielded by the median heuristic (bandwidths of 1, 5, and 10 for the three datasets, respectively, as shown in the main text) are quite near the peaks of the curves at which MMD most effectively distinguishes the distributions.
 Figure S7: Evaluating the effect of varying bandwidth parameters in Gaussian Kernel MMD for different sample sizes and dataset dimensionalities.
@@
 We show the mean and standard deviation over five runs of randomly sampled data.
-(a) $MMD^{2}$ distance with varying bandwidth parameters (0.5, 1, 2) for the 2d-MoG example shown in Fig.
+(a) $\mathsf{MMD}^{2}$ distance with varying bandwidth parameters (0.5, 1, 2) for the 2d-MoG example shown in Fig.
 1 compared to samples from a unimodal Gaussian approximation with the same mean and covariance.
-(b) $MMD^{2}$ distance with varying bandwidth parameters (1, 3, 5) for a ten dimensional standard normal distribution, for which the first dimension is shifted by one for the shifted example.
-(c) $MMD^{2}$ distance with varying bandwidth parameters (5, 10, 40) based on 10k samples from a standard normal distributions with varying dimensions (between 5 and 1000).
+(b) $\mathsf{MMD}^{2}$ distance with varying bandwidth parameters (1, 3, 5) for a ten dimensional standard normal distribution, for which the first dimension is shifted by one for the shifted example.
+(c) $\mathsf{MMD}^{2}$ distance with varying bandwidth parameters (5, 10, 40) based on 10k samples from a standard normal distributions with varying dimensions (between 5 and 1000).
 As in (b) the first dimension is shifted by one for the ’shifted’ dataset.
 Here we show one run due to computational costs.
-A.9 Details about scientific application examples
-For the motion discrimination task, we used the decision times of a single animal during both correct and erroneous trials with dot motion coherence of 12.8%, leading to a one-dimensional dataset of 587 samples.
-DDMs were implemented using the pyDDM toolbox (Shinn et al., 2020).
+We then choose a set of bandwidth parameters to compare across the scaling experiments of Section 3 (Fig.
+S7).
+Again, poor choices of bandwidth values give misleading results, but bandwidth choices guided by the median heuristic generally perform well.
+Figure S8: Comparison of Gaussian Kernel MMD to different MMD kernels without tunable parameters.
+We compare the performance of $\mathsf{MMD}^{2}$ as presented in Fig.8 with a Gaussian kernel with bandwidth parameters adjusted for each dataset (1,5,10) (top row) to an MMD implementation with a linear kernel (middle row, $\mathsf{MMD}_{lin}$; $k(x,y)=\langle x,y\rangle$) and an energy-distance based kernel, i.e., the kernel induced by the Euclidean distance Sejdinovic et al.
+(2013) (bottom row, $\mathsf{MMD}_{en}$; $k(x,y)=\|x\rVert^{p}+\|x\rVert^{p}-\|x-y\rVert^{p}$, with $p=2$).
+The experiments, parameters for the Gaussian kernel bandwidth (indicated in the y-labels), and sample sizes etc.
+are identical to Fig.
+8.
+Finally, we vary the kernel choice across the scaling experiments of Section 3 (Fig.
+S7), using both a linear kernel ($\mathsf{MMD}_{lin}$) and the distance-induced kernel corresponding to the standard energy distance ($\mathsf{MMD}_{en}$).
+As expected, the linear kernel fails to distinguish distributions with matching means (2d-MoG) but performs reasonably well for distributions with mean-offsets even at high dimensions.
+The energy kernel performs similarly to the Gaussian kernel, without the added dependence on sensitive hyperparameters.
+A.13 Additional results for ImageNet generative models
+We generated 50,000 images using an unconditional diffusion model explicitly trained on ImageNet 64x64, as well as 100,000 class-conditional images from a conditional variant of the same model (Dockhorn et al., 2022) (we refer to it as GENIE).
+For additional comparison, we also generated 50,000 images using a consistency model explicitly trained on ImageNet 64x64 (Song et al., 2023) (we refer to it as CM).
+We compare these generated images to the ImageNet 64x64 test set.
+For further comparison, we also evaluated image-generative models not specifically trained to reproduce ImageNet but designed for general-purpose image generation, such as Stable Diffusion and Midjourney (Rombach et al., 2022b, Midjourney, 2022).
+While these models might generate images that are more appealing to human observers, they do not necessarily produce images that align with the images contained in the ImageNet test set.
+We use the recently published million-scale dataset GenImage (Zhu et al., 2023).
+Especially, we include the models BigGAN (Brock et al., 2019), ablated diffusion model (ADM; Dhariwal & Nichol 2021), Glide (Nichol et al., 2022), Vector Quantized Diffusion Model (VQDM; Gu et al.
+2022), Wukong (Wukong, 2022), Stable diffusion 1.5 ( SD1.5; Rombach et al.
+2022b) and Midjourney  (Midjourney, 2022).
+It’s important to note that not all models are specifically trained to capture ImageNet 64x64 images.
+For instance, models like Stable Diffusion and Midjourney are trained on much larger datasets (Schuhmann et al., 2022, Lin et al., 2014).
+Additionally, most of the models mentioned, except for ADM and BigGAN, are text-to-image generative models.
+To minimize significant distribution shifts, these models were prompted with the phrase "photo of [ImageNet class]" Zhu et al.
+(2023).
+Furthermore, all the other models generate images of larger resolution, which we resized to 64x64.
+Hence, we also only compare low-resolution features of natural images.
+[…]
+Overall, this analysis highlights that the choice of metric (i.e., what features it compares) and the specific implementation details (such as the classifier in C2ST estimates) matter and can lead to varying results.
+Table S4: Evaluating discrepancy to the ImageNet test set.Each row presents various metrics computed on the Inception v3 embeddings of images.
+Columns correspond to different generative models.
+The initial three models are trained on ImageNet 64x64, serving as the reference point for comparison.
+Subsequent models are trained on alternative datasets or higher-resolution versions.
+In bold, we highlight the lowest value for each section of the table.
+GENIE CM BigGAN ADM Glide VQDM WK SD1.5 Midjourney
+FID $\mathbf{5.1\cdot 10^{0}}$ $5.3\cdot 10^{0}$ $1.2\cdot 10^{1}$ $1.1\cdot 10^{1}$ $1.1\cdot 10^{1}$ $\mathbf{9.8\cdot 10^{0}}$ $1.1\cdot 10^{1}$ $1.2\cdot 10^{1}$ $1.0\cdot 10^{1}$
+SWD $22.3\cdot 10^{-}$ 2 $\mathbf{2.2\cdot 10^{-}2}$ $25.3\cdot 10^{-}$ 2 $25.1\cdot 10^{-}$ 2 $24.9\cdot 10^{-}$ 2 $\mathbf{4.0\cdot 10^{-}2}$ $24.8\cdot 10^{-}$ 2 $25.0\cdot 10^{-}$ 2 $24.5\cdot 10^{-}$ 2
+[…]
+[…]
+[…]
+[…]
+[…]
+A.14 Details about scientific application examples
+For the motion discrimination task, we used the decision times of a single animal during both correct and erroneous trials with dot motion coherence of 12.8%, leading to a one-dimensional dataset of 1023 samples.
+From these 80% were used as a train set and 20% as a test set DDMs were implemented using the pyDDM toolbox (Shinn et al., 2020).
 DDM1 used a linear drift and exponential decision boundaries.
@@
 Both were sampled 1,000 times to create the two synthetic datasets.
-The real chest X-ray dataset consists of 70,153 samples, the generated datasets from PGGAN and SD consist of 10,000 and 2,352 samples respectively.
+The real chest X-ray dataset consists of 70,153 train samples and 25,596 test samples, the generated datasets from PGGAN and SD consist of 10,000 and 2,352 samples respectively.
 In both applications we computed metrics between pairs of 10 random subsets from the compared distributions (scatter points on the violin plots).
@@
 […]
+unsrtnat_abbr
 […]
+refs_table